A Complete Bibliography of Proceedings of the ACM on Computer Graphics and Interactive Techniques

(PACMCGIT)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

20 August 2024 Version 1.03

Title word cross-reference

3 [CL24, DAI^{+}18, FLS^{+}22, FXX^{+}22, LL23, NF21, PKK^{+}24, PPAM23, STSK20, WDM24]. k [KB24].

-DOP [KB24].

/Good [Rüs21].

60Hz [RW21].

À-Trous [DHD24]. Accelerate [WMZ^{+}20]. Accelerated [LVY^{+}20, SKRS24, KKSM18]. Accuracy [BBD24, Wan18]. Accurate [KK19]. acoustic [CASR21]. ACT2G [TWT^{+}23]. Action [NBM23]. Active [TB23]. Active-Set [TB23]. Adaptation [NBD^{+}24]. Adapted [DWBV24]. Adaptive [EK18, MYS20, MAK23, RHH^{+}22, SPD18, WSG19, XO21, YZK^{+}19, AÖG24, XOKN20]. Address [SLY20]. Advances [Hes22]. Advection [NZT19]. Advection-Reflection [NZT19]. Adversarial [SPT24, Two22, YKK^{+}23]. African [OBC22]. After [PWSF24]. Age [OSM^{+}22]. Agencies [DCB23]. Agency [HGR^{+}22]. Agent [HG19, Wei23, YW19, YKK^{+}23]. Agent-based [YW19]. agents [CASR21]. Agnostic [DB24]. AI [CASR21, HHYG24, Mes24, NBM23, Oh23, SG22, Two23, YEK24, ZCL24]. AI-aided [SG22]. AI-driven [HHYG24]. AI-drone [Oh23]. aided [SG22]. Aiming [KKS^{+}20]. Algorist [Zho24]. Algorithm [DB24, WH21]. Algorithm-Agnostic [DB24]. Algorithms [DXS21, HRZ^{+}19, Zho24]. Aliasing [CLS^{+}21, FHT^{+}24]. Alpha [Yuk18]. Altered [WH21]. Alternating [ANAM^{+}20]. Am [RW21]. Ambient [VSE21, SBE22]. Analysis [SK23]. Ancient [EZ22]. Andean [MD21]. Angle [WK23]. Angle-Based [WK23]. Angular [RCH^{+}23]. Animals [YBB^{+}21]. Animatable [SNL23]. Animated [SLR^{+}22]. Animation [FLS^{+}22, FG24, HYW^{+}23, LFFJ^{+}23, YY21, YW19]. Animations [JWLC23]. Anisotropic [KHDN22]. Annotations [AHAC^{+}24]. Anti [CLS^{+}21]. Anti-Aliasing [CLS^{+}21]. Anything [MBN^{+}24]. aperiodic [LSG24]. Apparent [LD18, RPP21]. Appearance [ANEK21, GCKM24]. Application [MDD^{+}22, Dup20]. Applications [CKY^{+}22]. Approach [HLX^{+}23, NBD^{+}24, RPP21, STSK20, XK21]. Approaches [FJdNJ22]. Approximate [RPHD20]. Approximation [LPW20, TYS20, YEK24]. Approximations [STK20]. Aquarium [LA22]. Aquaterrestrial [NBM23]. Arabic [SPT24]. Arbitrary [BD22, BBM24]. Archaeology [Hol22]. Architecture [JWB^{+}21]. Arm [Two22]. Art [Fuj21, HHYG24, ZRL21, SG22, ZCL24]. Artificial [CG23, LA22]. Artist [FG24]. Artistic [Kur24, NF21, SPT24]. Assemblage [Mes24]. Assisted [MDD^{+}22]. Assistive [KLXvdP21]. Athletes [SWF^{+}22]. Attention [TWT^{+}23]. Attention-based [TWT^{+}23]. Attribute [PKM22]. Audiences [BLLJ^{+}24]. Audio [FLS^{+}22, LH23]. Audio-Text [FLS^{+}22]. Auditory [HK24]. Augmented [LHK^{+}24].

Autonomous [BYH23]. Auxiliary [HOKY23]. Avatars [CO18, ZM23]. Averaging [AS22, DXS21]. Aware [WDM24, AHAC^{+}24, LHL24]. Awareness [LHK^{+}24].

B [ZWWL23]. B-Spline [ZWWL23]. Bad [Rüs21]. Ball [ZWWL23]. Based [DHD24, GW23, HCK23, JDZD19, JFS^{+}21, KJM22, LBR^{+}18, LBG18, LHR^{+}23, LFFJ^{+}23, MCC22, MKKP18, NF24, PJL21, RPHD20, RPP21, SHS^{+}21, SK23, SWF^{+}22, SyC23, SDY^{+}18, TKC21, WGD^{+}23, WK23, WU23, XK21, Yuk24, AÖG24, ARM^{+}19, BWL18, FJdNJ22, HYW^{+}23, KK19, Muk18, PLRD21, RWY^{+}23, TWT^{+}23, TTK^{+}21, Wei23, WH21, YW19, YY21, YKK^{+}23]. Beauty [DCB23]. Behavior [MDD^{+}22]. Behaviour [NBM23]. Below [Hes22]. Better [AS22, RW21]. Between [Two22]. Betweening [SSKS23]. Bidirectional [NF24]. Billion [SKW22]. Binary [BD24, Dup20]. Biomimetic [Oh23]. Biosignals [HCK23, YBB^{+}21]. Bird [Oh23]. bisection [Dup20]. Blend [PKM22]. Blending [HN18, ZWWL23, LSG24]. Blue [LWLY22]. Blur [OCW22]. Blurred [CL24]. Bodies [MWW18, NF21]. Body [LAJB18, OSM^{+}22, ZM23]. Bodylab [ZM23]. Bones [Muk18]. Botanical [QLZF21]. Boundary [DB23, YIFO24]. Boundary-Respecting [DB23]. Bounded [TE24]. Bounding [BDTD22, BMB^{+}24]. Box [TB23, XTS^{+}23, YH23, YW19]. Box-Constrained [TB23]. box-manipulation [YW19]. Brain [KBM23]. Brain-Computer [KBM23]. BRDF [TE24]. Bringing [KHDN22]. Buffers [HCNS18]. Builder [VKJ^{+}18]. Building [KK19]. Bunny [YIFO24]. BVH [VKJ^{+}18]. By-Example [HN18].

C [SFPO22]. Cache [KKI^{+}18]. Caches [TWS^{+}24]. Caching [KDSD24, SKRS24, TWS^{+}24]. Calculate [Wan18]. Calligraphic [SPT24]. Camera [CL24, SLR^{+}22]. Can [Far21]. Cangjie [ZRL21]. Caps [PD19]. Capture [BYH23, SPL^{+}21]. Carlo [DB24, HOKY23]. Carol [GCY^{+}22]. Cars [RPHD20]. Caustics [GHH24]. Cell [LHK^{+}24]. centric [HRZ^{+}19]. Ceramics [EZ22]. Change [NBM23]. Character [GCKM24, XK21, YY21]. Characters [TTC24, YKK^{+}23]. Cheat [JFS^{+}21]. Child [Two22]. Chinese [CYS23, Zho24]. Christmas [GCY^{+}22]. Church [HK24]. Classical [Zho24]. Classification [SWF^{+}22]. Climate [NBM23, RN24b]. Clouds [GSNK24, LL23, SHW24]. Clustering [BDTD22, BMB^{+}24]. Co [GCKM24, SFPO22, Two22]. Co-opting [SFPO22]. Co-robotic [Two22]. Co-solving [GCKM24]. codec [KKSM18]. Coding [PKM22]. Cognitive [Mes24]. Coiled [SWP^{+}23]. Collaborating [GCKM24]. Collaborative [Two22]. Collision [MEM^{+}20, SK23, TLTM18]. Communication [HG19]. Communing [Two23]. Compacted [SLY20]. Companion [YBB^{+}21]. Comparative [SLW^{+}21]. comparison [WJB23]. Compensated [SGH18]. Competitive [JFS^{+}21, Yuk24]. Compliant [SyC23]. Complicated [JWB^{+}21]. Components [BD24]. Compositional [AHAC^{+}24]. Compressing [BBM24]. Compression [LMSS18, PKM22, SLY20, vdLSE20]. Computer [KBM23, Sim21]. Concert [HK24]. Concurrent [BD24, Dup20, WGD^{+}23]. Conditions [KKS^{+}20]. Conductors [MD21]. Conflict [Cha24]. Confronting [BLLJ^{+}24]. Conical [XJZ^{+}22]. Conserving [RCH^{+}23]. consistent [SBE22, WJB23]. Constitutive [SLX^{+}23]. Constrained [RHH^{+}22, TB23]. Constraint [SyC23]. Construction [BDTD22, BMB^{+}24]. Contact [HGG^{+}19, LB19, RCH^{+}23]. Contacts [BBGB20]. Content [YZK^{+}19]. Context [LHL24, WDM24, SG22]. Context-aware [LHL24]. Contextualizing [HK24]. Continuity [MD21]. Contrastive [TWT^{+}23]. Control [BWL18, KBFF^{+}21, XO21, XTS^{+}23, XXA^{+}23, XK21]. Converging [DB24]. Conversations [JDZD19]. Conversion [Muk18]. Convolution [JWB^{+}21, TKC21]. Convolutional [CAS22]. Cooperative [SSH^{+}21, YW19]. Coordinating [HG19]. Corotated [KBFF^{+}21]. Correction [ERHR22, WJB23]. Cosines [KHDN22]. Cost [MUEM22]. Coupling [ANEK21, WU23]. CPU [SLY20]. CPU/GPU [SLY20]. Craft [EZ24]. Craft- [EZ24]. Crafting [MXZ22]. Crash [CG23]. Create [FXX^{+}22, KBM23]. Creating [ZCL24]. Creation [LMP22]. Creative [Hes22, Two23, SG22]. Critical [RN24b]. Crowd [DMH^{+}21, TZW24]. Crowds [TZW24]. CSG [ZCL18]. Cubic [TYS20]. Culling [KSK24, TLTM18, XJZ^{+}22, YH23]. Curl [DB23]. Curl-Noise [DB23]. Curls [SWP^{+}23]. Curve [ZWWL23]. Curves [TYS20]. Cyclic [JWLC23]. Cymatics [HCK23].

D [CL24, DAI^{+}18, FLS^{+}22, FXX^{+}22, LL23, NF21, PKK^{+}24, PPAM23, STSK20, WDM24]. D-Aware [WDM24]. Dance [XBN^{+}23, PW23, PW23]. dance-musical [PW23]. Dances [OBC22]. Data [BDBS22, Day21, HMES20, KK19, OSM^{+}22, SLY20, TKC21, WP22]. Data-Driven [BDBS22, TKC21]. DCGrid [RHH^{+}22]. Deblur [CL24]. Deblur-GS [CL24]. Decompression [GSNK24]. Deep [BBD24, CWZ^{+}18, DAI^{+}18, FJdNJ22, JDZD19, LWM19, NBD^{+}24, OBC22, SDY^{+}18, TKC21]. Deferred [MYS20]. Deformable [CDGB19, LWM19]. Deformation [LSL^{+}24, WK23]. Deformer [BBGB20]. DEM-based [AÖG24]. Democratic [Cha24]. demonstrations [YKK^{+}23]. Denoisers [DHD24]. Denoising [DB24, HOKY23, HMES20, HHM23, TLP^{+}22]. Dense [BBM24]. Dependent [LBG18]. deposition [RDBC24]. Depth [FHSS18, HCNS18, VSE21, KKSM18]. derivatives [WJB23]. Design [BWL18, Far21, MDD^{+}22, MWW18, NBM23]. Designed [FJdNJ22]. Desire [CG23]. Destruction [WFM21]. Detailed [RN24a]. Detection [FHT^{+}24, JFS^{+}21]. DGF [BBM24]. Difference [ANAM^{+}20]. Different [CYH^{+}24]. Differentiable [DHB24, DB23, SyC23, SLX^{+}23]. Diffuse [LHR^{+}23, dDB22]. Diffusion [CG23]. DiffXPBD [SyC23]. Digital [BLLJ^{+}24, EZ22, HLX^{+}23, HCR^{+}21, Hes22, NF21, OSM^{+}22, YIFO24, EZ24]. Dimension [GHH24]. Dimensional [CDGB19, QLZF21]. Dimensions [SPT24]. Disconnect [Day21]. Discontinuities [DB23]. Discrete [MAK23]. Disk [WS18]. Displacement [LMSS18]. Display [GW24]. Displays [FSFS23, LD18]. Dissecting [OSM^{+}22]. Distance [MEM^{+}20, YEK24]. Distribution [Yuk18]. Divergence [Wan18]. Diverse [HGG^{+}19, PJL21]. Domain [NBD^{+}24]. Domains [PJL21]. DOP [KB24]. Doppelgänger [GCKM24]. Doppelgängers [BLLJ^{+}24]. DP [RN24a]. Drawing [TTC24, Two22]. Drawn [FG24]. Dream [SG22]. Driven [BDBS22, FLS^{+}22, LJC24, TKC21, HHYG24]. Drone [Oh23]. drop [Fuj21]. Dual [LVY^{+}20]. Dual-Split [LVY^{+}20]. Ductile [WDG^{+}19]. Dune [TK23]. During [HGR^{+}22, GCKM24]. Dynamic [GW23, JWB^{+}21, LHR^{+}23, MLD^{+}24]. Dynamics [CO18, SyC23, TZW24, Wei23, WU23].

Easily [FXX^{+}22]. EasyVRModeling [FXX^{+}22]. edge [Dup20]. Editing [JKK^{+}23, WDM24]. Effect [WSX^{+}22]. Effects [GW23, GCKM24, RW21, WSG19]. Efficient [AHGB19, CASR21, DXS21, FSFS23, LSL^{+}24, MYS20, MLD^{+}24, NF24, SCNW19, SGH18, Tok23, YY21, YEK24]. Eigenanalysis [WK23]. ejections [LRC^{+}22]. Elastic [ARM^{+}19, BBGB20, CDGB19, KBFF^{+}21, MWW18, RCH^{+}23]. Elastic-rigid [BBGB20]. Elasticity [LFFJ^{+}23]. Electronic [MD21]. Elucidated [Sim21]. Embodied [BDWG21, PW23]. Embodiment [ECM23]. Endeavors [Cha24]. Energies [SK23, WK23]. Energized [LAJB18]. Energy [KBFF^{+}21, SGH18]. Energy-Compensated [SGH18]. Engaging [KBM23]. Enhancement [LD18]. Environment [HCK23, WH21]. Environmental [XJZ^{+}22]. Environmentalist [NBM23]. Ephemera [HHYG24]. Erratics [LMP22]. Error [CKY^{+}22]. Estimation [DHB24, GHH24, KJM22, SPD18]. Evaluating [Wan18, XXA^{+}23]. Evaluation [DMH^{+}21]. Evaluator [ANAM^{+}20]. Event [GHH24]. Events [FJdNJ22]. Evolutionarily [PPAM23]. Evolutionarily-trained [PPAM23]. Example [HN18, TTK^{+}21]. Example-based [TTK^{+}21]. Expansion [LU24]. Experience [HCR^{+}21, WSX^{+}22, WSG19, ZRL21, CYS23]. Experiencing [BYH23]. Experimental [KBM23]. Explicit [KDSD24]. Exploration [RN24b]. Explorations [DCB23]. Explore [CG23]. Exploring [SG22, PW23]. Expression [MXZ22]. Expressive [EZ22, FLS^{+}22]. Extracted [Wan18]. Eyck [Sim21]. Eye [BBD24, BDWG21, GW23, JDZD19, KWK22, MBN^{+}24, NBD^{+}24, Oh23, SWF^{+}22, TTC24]. Eye-Based [GW23]. Eye-gaze [TTC24]. Eye-Tracking [BBD24, BDWG21, NBD^{+}24]. fabrication [EHZ23]. Face [KK19]. FaceBlit [TTK^{+}21]. FaceFolds [MLD^{+}24]. Faces [MUEM22, MLD^{+}24]. FaceType [MXZ22]. Facial [FLS^{+}22, TTK^{+}21]. Fast [DHD24, KBFF^{+}21, LSG24, VKJ^{+}18, WS18, Wan18, WDM24, Wei23, DWBV24]. faster [GBW20]. Feature [MSK22]. Feature-Wise [MSK22]. Features [HOKY23, MBN^{+}24, SFPO22]. Feedback [MCC22]. Feminism [Far21]. fidelity [HLX^{+}23]. Field [FSFS23, FHSS18, GW24, MEM^{+}20, SLMM22, KKSM18]. Fields [CWZ^{+}18, JKK^{+}23, YEK24]. fighting [YKK^{+}23]. Figures [NF21]. filmic [PW23]. Filmmaking [KBM23]. Filter [DWBV24, TY18]. Filter-Adapted [DWBV24]. Filtering [AS22, CLS^{+}21, CAS22, OCW22, PWSF24, SPD18]. Filters [STK20]. Finding [Yuk22]. Fine [EHZ23]. first [WJB23]. first-order [WJB23]. Five [Two22]. Five-Year-Old [Two22]. Fixed [AS22]. Fixed-Point [AS22]. Flexible [KLXvdP21]. FLIP [ANAM^{+}20]. Floating [YIFO24]. Flow [CYH^{+}24, RPHD20, RPP21, WXCT19]. Flows [DB23, WFM21]. Fluid [NF24, RHH^{+}22, TY18, WXCT19, YCL^{+}19]. fly [KKT^{+}18]. Footage [FJdNJ22]. Footprint [PKK^{+}24]. Forces [KLXvdP21]. Format [BBM24]. Foveated [MDZV18]. Fracture [LAJB18, WDG^{+}19]. Framework [HRZ^{+}19, LSL^{+}24, MKVH24, PPAM23]. Free [PLRD21]. Free-view [PLRD21]. Freeze [PW23]. Friction [ANEK21]. Frictional [HGG^{+}19]. Friendly [BBM24, LU24]. Frustum [SKRS24]. full [ZM23]. full-body [ZM23]. functional [MMMC21]. Functions [Kir18]. Fundamental [XXA^{+}23].

G [RW21]. G-SYNC [RW21]. G2 [ZWWL23]. Game [BD24, FJdNJ22, SLW^{+}21, SSH^{+}21, YEK24]. Gameplay [FJdNJ22, RW21, WSX^{+}22, WSG19]. Games [CYH^{+}24, WSX^{+}22, YZK^{+}19, Yuk24]. Gaming [JFS^{+}21]. GAN [AHAC^{+}24, WXCT19, XK21]. GAN-Like [XK21]. Garments [LSL^{+}24]. Gaussian [CL24, PKK^{+}24]. Gay [Zho24]. Gaze [AHAC^{+}24, KJM22, KWK22, TTC24]. Gaze-aware [AHAC^{+}24]. General [HRZ^{+}19]. Generalized [Oga20, SLX^{+}23]. generate [Tri24]. generated [EZ24]. Generation [JDZD19, KOF^{+}24, LBG18, SHW24, SDY^{+}18, TWT^{+}23, WS18]. Generative [PJL21, SPT24, Two22, EHZ23]. Genetic [WH21]. Genetic-algorithm-based [WH21]. Geometric [AHGB19, CLS^{+}21, WJG^{+}21]. Geometry [BBM24, KKT^{+}18, QLZF21, vdLSE20]. Gesture [TWT^{+}23]. GGX [TE24, KHDN22]. Ghost [EZ22]. GigaVoxels [RN24a]. Glint [CLS^{+}21]. Global [LHR^{+}23, SPL^{+}21, TWS^{+}24]. Good [Rüs21]. GPT [Mes24]. GPT-ME [Mes24]. GPU [DHD24, HYW^{+}23, KKI^{+}18, KKSM18, KOF^{+}24, LU24, RHH^{+}22, SLY20]. GPU-accelerated [KKSM18]. GPU-friendly [LU24]. GPUs [TLTM18]. Gradient [DHB24, SPD18]. Graph [PJL21]. Graph-Based [PJL21]. Graphics [Yuk22]. Graphs [KOF^{+}24]. Gravity [Kur24]. Grid [LY19, RHH^{+}22, RCH^{+}23]. Grids [NF24]. Group [Wei23]. GROUPTHINK [HGR^{+}22]. GS [CL24]. Guidance [HOKY23]. Guided [KSK24, KW21, OCW22, XOKN20].

H [BMB^{+}24]. H-PLOC [BMB^{+}24]. Hair [CAS22, HYW^{+}23, RL18, SWP^{+}23]. Half [YIFO24]. Half-Digital [YIFO24]. Half-Physical [YIFO24]. Halfedge [BD22]. Hand [FG24]. Hand-Drawn [FG24]. Handcrafting [Kur24]. Haptic [MCC22]. Hardware [BBM24, LVY^{+}20, MYS20, VKJ^{+}18, WMZ^{+}20]. Hardware-Accelerated [LVY^{+}20]. Hardware-Friendly [BBM24]. Harmonics [RSS^{+}24]. Hashing [TLTM18, ZCL18]. HCSG [ZCL18]. HDHumans [HLX^{+}23]. Head [JDZD19, LD18]. Head-mounted [LD18]. Helper [Muk18]. Heritage [Hes22]. Heterogeneous [PHM^{+}21]. HeterSkinNet [PHM^{+}21]. Hierarchical [BMB^{+}24, HMN18, XTS^{+}23]. Hierarchies [KB24, WP22]. Hierarchy [BDTD22, BMB^{+}24, LY19]. High [HLX^{+}23, HN18, KKS^{+}20, LMSS18, STK20, VKJ^{+}18, Yuk22, vdLSE20, KKSM18]. High-fidelity [HLX^{+}23]. High-Performance [HN18, STK20, Yuk22]. High-Quality [VKJ^{+}18, KKSM18]. High-Resolution [LMSS18]. HIP [MKVH24]. HIPRT [MKVH24]. Histogram [HN18]. Histogram-Preserving [HN18]. History [HK24]. Htex [BD22]. Human [MUEM22, MSK22, Mes24, NF21, ZRL21]. Human-Machine [ZRL21]. Humans [HLX^{+}23]. Hybrid [DPM24, HLX^{+}23, HGG^{+}19, HYW^{+}23, KSK24, PLRD21, RCH^{+}23]. Hyperparameter [YY21].

Identify [FJdNJ22]. Illumination [LHR^{+}23, TWS^{+}24]. Image [Hol22, LBR^{+}18, PLRD21, RPHD20, STK20, WDM24, SG22]. Image-Based [LBR^{+}18, RPHD20, PLRD21]. Images [ANAM^{+}20, CL24, FHT^{+}24, EZ24]. Imaginary [LZY21]. Imitation [XK21, YKK^{+}23]. Immersive [CYS23, FXX^{+}22, HCR^{+}21, HHYG24]. Impact [CYH^{+}24]. Implicit [KBFF^{+}21]. Importance [EK18, LHR^{+}23, XOKN20]. Importance-Based [LHR^{+}23]. Impressions [MXZ22]. Improves [KKS^{+}20]. Improving [NBD^{+}24]. In-Betweening [SSKS23]. In-Depth [HCNS18]. Incompressible [DB23]. Increase [BBD24, RPP21]. Independent [MKKP18, BBGB20, LHL24]. Indigenous [MD21]. Indoor [JWB^{+}21]. Information [WJG^{+}21]. Input [KKS^{+}20]. Inputs [RWY^{+}23]. Insects [WH21]. Installation [Fuj21, HHYG24, SLMM22, PW23, ZCL24]. Instant [SLR^{+}22, TTK^{+}21]. Instruments [BDWG21]. Interaction [HCK23, SHS^{+}21, YKK^{+}23]. Interactive [CO18, Fuj21, GHH24, HHCM21, HHYG24, JKK^{+}23, LRC^{+}22, MCC22, MWW18, PPAM23, SPL^{+}21, SLMM22, SNL23, SDY^{+}18, XK21, ZRL21, PW23, SG22, ZCL24]. Intercorporeality [LHK^{+}24]. Interface [BDWG21, KBM23]. Interfaces [dDB22]. Intersection [BYH23, YH23]. Intersections [Res22, WMZ^{+}20]. Intersector [RL18]. Interval [Tri24]. intervals [Tri24]. Intervention [Kur24]. Invariant [ARM^{+}19]. Inverse [SLX^{+}23]. Investigating [FJdNJ22]. Invisible [Rüs21]. Island [HK24]. Isosurface [Wan18]. Iteration [RCH^{+}23].

Jan [Sim21]. Jigsaw [GCKM24]. Joint [FLS^{+}22, HHM23, TLP^{+}22]. juggling [LH23]. Julian [Day21].

Kernel [MDZV18]. Kestrel [Oh23]. KineCAM [SLR^{+}22]. Kiss [CG23]. Kiss/Crash [CG23]. Know [OSM^{+}22].

Labels [RPHD20]. Labor [Rüs21]. Language [HHYG24, Tem24]. Large [BD24, RN24a]. Large-Scale [BD24]. Late [KKS^{+}20]. Latency [KKS^{+}20]. Layer [FHSS18]. Layered [dDB22]. Layouts [JWB^{+}21]. Learned [WJG^{+}21]. Learning [AHAC^{+}24, BBD24, CO18, FHT^{+}24, HG19, JDZD19, LSL^{+}24, OBC22, RN24b, RPP21, SLX^{+}23, TZW24, TWT^{+}23, TKC21, XK21, YW19]. Learning-Based [JDZD19]. Leaves [SLMM22]. Less [RN24a]. Level [NF24, STSK20]. Life [LA22]. Lifted [SWP^{+}23]. Light [CWZ^{+}18, ERHR22, FSFS23, GW24, KSK24, Oga20, KKSM18]. Lightcuts [LY20]. Lighting [DPM24, LWLY22, LY19, XJZ^{+}22]. Lightning [FG24]. Lights [EK18, LPW20, PW23]. Lightweight [WDM24]. Like [XK21]. Limited [AHAC^{+}24]. Linear [RCH^{+}23, WJB23]. Linearly [KHDN22]. Liquid [TKC21]. Liquids [AHGB19, RPP21]. Live [HGR^{+}22]. Living [WH21]. Local [MEM^{+}20, MSK22]. Locally [BDTD22, BMB^{+}24]. Locally-Ordered [BDTD22, BMB^{+}24]. Loco [XTS^{+}23]. Loco-Manipulation [XTS^{+}23]. Locomotion [MSK22, WGD^{+}23]. LOD [LBG18, SHW24]. Long [WMZ^{+}20]. longest [Dup20]. Lossily [BBM24]. Lossless [YZK^{+}19]. Lossy [vdLSE20]. Low [MUEM22]. Low-Cost [MUEM22].

MAAIP [YKK^{+}23]. Machine [DCB23, EZ22, RN24b, ZRL21]. Machine-Microbial [DCB23]. Machines [LA22]. Making [NF21]. Mandarin [TTC24]. Manifold [CDGB19, GHH24, LJC24]. Manifolds [MLD^{+}24, SSKS23]. ManiLoco [WGD^{+}23]. Manipulation [WGD^{+}23, XTS^{+}23, YW19]. Many [EK18]. Map [CLS^{+}21, SDY^{+}18]. Mapping [Fuj21, STSK20]. Mappings [FSFS23]. Mask [Far21]. masks [GBW20]. Massive [GSNK24]. Massively [KKT^{+}18]. Massively-Parallel [KKT^{+}18]. Matchmaking [Yuk24]. Material [HGG^{+}19, WDG^{+}19]. Materialization [EZ22, EZ24]. Materials [HGG^{+}19, dDB22]. Mathematical [MFKK24]. ME [Mes24]. Measurement [LZY21]. Media [Hol22]. Medical [HMES20]. Meditation [HCR^{+}21]. Memories [YBB^{+}21, ZCL24]. Memory [PKK^{+}24, RHH^{+}22]. Memory-Constrained [RHH^{+}22].

Merging [Zho24]. Mesh [BD22, NF24, Tri24, WS18]. Meshed [MLD^{+}24]. Meshlets [BBM24]. Method [HGG^{+}19, HYW^{+}23, WGD^{+}23, Wan18, WDG^{+}19]. Methods [SLW^{+}21, WJB23]. Metric [DMH^{+}21, LBG18, Wan18]. Microbial [DCB23]. Micromaps [WD24]. Micropolar [LFFJ^{+}23]. Mimicry [WH21]. Mind [KWK22]. Ming [HCR^{+}21]. Misinformation [RN24b]. Mitigation [SLW^{+}21]. Mixed [CDGB19, CNI^{+}20, SSH^{+}21]. Mixed-Dimensional [CDGB19]. Mixed-Reality [SSH^{+}21]. Mobile [CNI^{+}20, KWK22]. Mode [KBFF^{+}21]. Model [CDGB19, FLS^{+}22, JDZD19, KJM22, MBN^{+}24, PJL21, SWP^{+}23, SLX^{+}23]. Model-Based [KJM22]. Modeling [LB19, NF21, ZCL18]. Modelling [MSK22]. Models [CG23, FXX^{+}22, KK19, MFKK24]. Modern [HK24, KKI^{+}18]. Modes [CYH^{+}24]. Modified [SLY20]. Modulated [KLXvdP21]. Moment [MKKP18]. Moment-Based [MKKP18]. Momentum [RCH^{+}23]. Money [SLMM22]. Monte [DB24, HOKY23]. Mother [Rüs21]. Motion [CL24, JDZD19, KLXvdP21, MSK22, OCW22, PJL21, PPAM23, RWY^{+}23, SPL^{+}21, SCNW19, SSKS23, TZW24, XXA^{+}23, YZK^{+}19]. Motor [EHZ23]. mounted [LD18]. Movement [LZY21, XBN^{+}23]. MPM [SLX^{+}23]. Multi [DAI^{+}18, FHT^{+}24, FHSS18, HG19, LB19, LL23, Wei23, WXCT19, YKK^{+}23]. Multi-Agent [HG19, Wei23, YKK^{+}23]. Multi-Layer [FHSS18]. Multi-part [LL23]. Multi-Pass [WXCT19]. Multi-Resolution [LB19]. Multi-Task [FHT^{+}24]. Multi-View [DAI^{+}18]. Multigrid [AHGB19]. Multilevel [LSL^{+}24, TB23]. Multiple [PJL21]. Multiview [GW23]. Musical [RN24b, PW23]. My [OSM^{+}22, GCKM24].

Narrow [TY18]. Narrow-Range [TY18]. Nature [WH21]. Navigation [HG19]. NeRF [SKRS24, WDM24]. NeRFahedron [SNL23]. NeRFshop [JKK^{+}23]. Network [PHM^{+}21, Two22]. Networks [CAS22, ECM23, FJdNJ22, JWB^{+}21, PPAM23, SCNW19, SPT24, SDY^{+}18]. Networks-based [FJdNJ22]. Neural [CAS22, ECM23, FJdNJ22, HMES20, HHM23, JKK^{+}23, NBD^{+}24, PPAM23, RPP21, SNL23, SCNW19, SDY^{+}18, TLP^{+}22, YEK24]. NeuroDog [ECM23]. Next [GHH24]. Next-Event [GHH24]. Nine [Zho24]. Noise [Day21, DB23, HN18, Kir18, LWLY22]. Non [CDGB19, DHB24, Kir18, Muk18]. Non-Differentiable [DHB24]. Non-Manifold [CDGB19]. Non-periodic [Kir18]. Non-rigid [Muk18]. Nonlinear [CO18]. Normal [CLS^{+}21, SDY^{+}18]. Novel [HRZ^{+}19].

Object [WP22, WGD^{+}23]. Objects [CDGB19, LWM19, GBW20]. obstacles [RDBC24]. Occlusion [VSE21, SBE22]. ocean [LSG24]. Oculomotor [MFKK24]. Old [Two22]. Oliver [Day21]. Olympus [Tem24]. Omnidirectional [YEK24]. On-Surface [TWS^{+}24]. On-the-fly [KKT^{+}18]. One [DHB24]. Opacity [WD24, GBW20]. Operating [EHZ23]. Operator [HN18]. Optimization [KLXvdP21, MEM^{+}20, YY21]. Optimized [KB24, LBR^{+}18]. Optimizing [KDSD24, KKI^{+}18]. opting [SFPO22]. Order [MKKP18, NZT19, WJB23]. Order-Independent [MKKP18]. Ordered [BDTD22, BMB^{+}24]. orientable [LSG24]. Ornamentation [MDD^{+}22].

Paglen [Day21]. Painter [SG22]. painting [ZM23]. paintings [CYS23]. Paradigm [BDBS22]. Parallel [AÖG24, BDTD22, BMB^{+}24, HRZ^{+}19, KKT^{+}18, TLTM18, WP22]. Parameters [MFKK24]. part [LL23]. Particle [NF24, RCH^{+}23, RPP21, TKC21]. Particle-Based [NF24, RPP21, TKC21]. Particle/Grid [RCH^{+}23]. party [JDZD19]. Pass [WXCT19, XOKN20]. Passthrough [CNI^{+}20]. Patch [LBG18]. Patch-Based [LBG18]. Path [HMES20, HHCM21, KDSD24, KSK24, WP22, WSD24]. Pattern [BWL18]. patterns [RDBC24]. PDF [Tok23]. Peace [HK24]. Penalty [SK23]. Penalty-Based [SK23]. Per-Halfedge [BD22]. Per-Subject [MFKK24]. perception [CASR21]. Perceptions [DPM24, YIFO24]. Perceptually [DMH^{+}21]. Perceptually-Validated [DMH^{+}21]. Performance [FJdNJ22, HN18, HGR^{+}22, LHL24, STK20, WSX^{+}22, WSG19, Yuk22, LH23]. performing [ZM23]. periodic [Kir18]. Permutation [PKM22]. Perspectival [Sim21]. Perturbations [WJG^{+}21]. Phantom [RL18]. Phase [SSKS23]. Phases [MSK22]. Photographs [SLR^{+}22]. Photometric [LPW20]. Photon [LJC24, SGH18]. Photon-Driven [LJC24]. Physical [JWLC23, YIFO24]. Physically [LA22, LFFJ^{+}23]. Physically-Based [LFFJ^{+}23]. Physics [HYW^{+}23, KK19, MCC22, RWY^{+}23, SLX^{+}23, XK21, YY21, YKK^{+}23]. Physics-Based [MCC22, XK21, HYW^{+}23, KK19, RWY^{+}23, YY21, YKK^{+}23]. Pier [LMP22]. Pipeline [FG24]. Pixel [HOKY23, TKC21]. Pixel-wise [HOKY23]. Planning [XTS^{+}23]. Plant [HCK23, MFKK24]. Plant-Environment [HCK23]. Plants [HCK23]. Plastic [Hol22]. Platform [LHL24]. Platform-independent [LHL24]. Player [SSH^{+}21, Yuk24]. Playing [RW21]. PLOC [BDTD22, BMB^{+}24]. PLOCTree [VKJ^{+}18]. plume [LRC^{+}22]. poems [CYS23]. Poetics [Rüs21]. Poetry [ZRL21].

Point [AS22, GW23, GW24, GSNK24, HGG^{+}19, LL23, SHW24, WDG^{+}19]. Points [SKW22]. Poisson [TB23, WS18]. Poisson-Disk [WS18]. Policies [XXA^{+}23]. Politics [Day21, Rüs21]. Polygonal [LPW20]. Polynomial [Yuk22]. Portals [Oga20]. Position [ARM^{+}19, SPL^{+}21, SyC23, Wei23, WU23]. Position-Based [SyC23, ARM^{+}19, Wei23]. possibilities [SG22]. Post [KKS^{+}20]. Post-Render [KKS^{+}20]. Precision [BBD24]. Precomputed [BDBS22, LWM19]. Preconditioner [TB23]. Predicting [CKY^{+}22]. Prediction [DAI^{+}18, PHM^{+}21, SPL^{+}21]. Predictor [LHL24]. Preservation [OBC22]. Preserving [HN18, LSL^{+}24, SLP21]. Pressure [TB23]. Prevention [Oh23]. Primitive [SNL23, WMZ^{+}20]. Priors [YKK^{+}23]. Probabilistic [EZ24]. Problems [DXS21, XXA^{+}23]. Procedural [DB23, Kir18, KOF^{+}24]. Processing [Hol22, KKT^{+}18, KKI^{+}18]. Production [RN24a]. Productions [GCY^{+}22]. Programming [Tem24]. Progressive [WS18]. Projected [PD19]. Projection [Fuj21, WS18]. Projection-Mapping [Fuj21]. Projective [FSFS23, KB18]. Proposition [Day21]. ProteusNeRF [WDM24]. PSCC [TLTM18]. Public [SLMM22]. Pupillary [ERHR22]. Pupillometry [ERHR22]. Puzzle [GCKM24]. pyroclastic [LRC^{+}22].

Quadratic [RSS^{+}24, TYS20]. Quadruped [ECM23]. Quality [DMH^{+}21, VKJ^{+}18, XBN^{+}23, KKSM18].

Racing [SLW^{+}21]. Radar [Hes22]. Radiance [BDBS22, JKK^{+}23, KDSD24, LWM19, MLD^{+}24, TWS^{+}24]. Range [TY18]. Rasterization [GSNK24, SKW22]. Rasterizer [DHB24]. Rates [CYH^{+}24]. Ray [KB24, LMSS18, LHR^{+}23, MKVH24, RL18, Res22, WMZ^{+}20, XJZ^{+}22, YH23, GBW20]. Ray-Hair [RL18]. Ray-Traced [XJZ^{+}22]. Ray/Primitive [WMZ^{+}20]. Ray/Ribbon [Res22]. Re [EZ22]. Re-Materialization [EZ22]. Real [CKY^{+}22, CNI^{+}20, CLS^{+}21, CG23, CAS22, DWBV24, GSNK24, KSK24, KOF^{+}24, LD18, LY19, LY20, LPW20, MUEM22, MSK22, PD19, SPD18, SKW22, SCNW19, TWS^{+}24, TK23, TTK^{+}21, TLP^{+}22, WSD24, WH21, XOKN20, XO21, XJZ^{+}22, YCL^{+}19, KKSM18, ZCL18]. Real-Time [CKY^{+}22, CLS^{+}21, CAS22, DWBV24, GSNK24, KSK24, KOF^{+}24, LY19, LY20, MUEM22, MSK22, TWS^{+}24, TK23, TLP^{+}22, WSD24, XJZ^{+}22, YCL^{+}19, CNI^{+}20, LD18, LPW20, SPD18, SCNW19, TTK^{+}21, WH21, XOKN20, XO21, KKSM18, ZCL18]. Realistic [PPAM23]. Reality [BYH23, BBD24, CNI^{+}20, CYH^{+}24, DPM24, ERHR22, HMN18, OSM^{+}22, SLW^{+}21, SWF^{+}22, SSH^{+}21, WSX^{+}22, YIFO24, ZRL21]. Realized [LA22]. Realtime [HYW^{+}23]. Recall [YBB^{+}21]. ReCollection [ZCL24]. Recolonization [NBM23]. reconFIGURE [BLLJ^{+}24]. Reconstruction [HHCM21, NF24, QLZF21, KKSM18]. Recovering [WJG^{+}21]. Redirected [SHS^{+}21]. Reducing [PKK^{+}24]. Reduction [GHH24]. Reflection [NZT19, RPHD20]. Reflex [ERHR22]. Region [LBR^{+}18]. Reimagining [Oh23]. reinforcement [YW19]. Reinventing [Kur24]. Related [NBM23]. Reliability [MFKK24]. Relighting [MUEM22]. ReMember [YBB^{+}21]. Render [KKS^{+}20, RN24a, WSX^{+}22]. Rendered [FHT^{+}24]. Rendering [DB24, DWBV24, FSFS23, FHSS18, GW23, GW24, GHH24, LY19, MLD^{+}24, MDZV18, PLRD21, RPHD20, SHW24, SNL23, SKRS24, TY18, dDB22, Tri24]. Replate [CSN18]. Representation [LL23]. Representations [CG23]. Resampling [Tok23]. Research [Kur24]. Resolution [LD18, LB19, LMSS18, RPP21, WSX^{+}22, WXCT19, vdLSE20]. Respecting [DB23]. ReSTIR [WSD24]. Retargeting [RWY^{+}23]. Rethinking [KJM22]. Reuse [KKT^{+}18]. Reversal [Day21]. Revisited [BDTD22]. Revisiting [KKI^{+}18]. Ribbon [Res22]. Rig [Muk18]. Rigid [LAJB18, BBGB20, Muk18]. Rigidification [MAK23]. Road [LMP22]. Robotic [Cha24, Far21, EHZ23, Two22]. Robotics [BDWG21]. Robust [ERHR22, JFS^{+}21, MEM^{+}20]. Robustness [BBD24]. Rods [ARM^{+}19]. Root [Yuk22]. Rotation [DXS21].

SAH [KB24]. SAH-Optimized [KB24]. SAM [MBN^{+}24]. Sample [WS18]. Samples [WS18]. Sampling [DWBV24, EK18, KKS^{+}20, KW21, LJC24, Muk18, PD19, TE24, XO21, XOKN20]. Sampling-based [Muk18]. Sand [TK23, RDBC24]. Scale [BD24]. Scan [Hol22]. Scatter [MYS20]. Scattering [WSD24, XOKN20]. Scene [JWB^{+}21, LBR^{+}18, RPP21]. Scenes [vdLSE20]. Schedule [DHD24]. Scheme [LMSS18]. Screen [TY18, SBE22]. Screen-Space [TY18, SBE22]. Script [SPT24]. Sculpting [MCC22, ZM23]. Search [Hol22, PW23]. Second [NZT19]. Second-Order [NZT19]. Seeing [BYH23]. Seepage [WFM21]. Segment [MBN^{+}24]. Segmentation [MBN^{+}24]. Selection [LBG18]. Self [GCKM24, MDD^{+}22, TLTM18]. Self-Collision [TLTM18]. Self-Shaping [MDD^{+}22]. Self-similar [GCKM24]. Semantic [RPHD20, ZRL21]. Serendipitous [NF21]. Set [TB23]. Setup [MUEM22]. Shader [LHL24, SFPO22]. ShaderPerFormer [LHL24]. Shaders [NF24, Tri24]. Shading [MYS20, PWSF24, Tri24, YZK^{+}19]. Shadow [CG23]. Shadows [XJZ^{+}22]. Shan [HCR^{+}21]. Shape [LL23, MWW18]. Shapes [LB19]. Shaping [MDD^{+}22]. Shells [MAK23]. Shot [MBN^{+}24]. Sickness [SLW^{+}21, WSX^{+}22]. Signed [MEM^{+}20]. Sim2Real [NBD^{+}24]. similar [GCKM24]. Similarity [Tok23]. SimLOD [SHW24]. Simplicial [CDGB19]. Simplification [LBR^{+}18]. Simulatable [QLZF21]. Simulation [LA22, MDD^{+}22, RHH^{+}22, SLP21, SWP^{+}23, SyC23, SLX^{+}23, TK23, TKC21, WDG^{+}19, WFM21, YCL^{+}19, AÖG24, LRC^{+}22, RDBC24]. Simulation-Assisted [MDD^{+}22]. Simulator [WSX^{+}22]. Simultaneous [SHW24]. single [XOKN20]. ’Sisyphus’ [Cha24]. skeleton [MMMC21]. Sketch [SDY^{+}18]. Sketch-Based [SDY^{+}18]. Sketching [DAI^{+}18]. Skill [Yuk24]. Skill-Based [Yuk24]. Skills [EHZ23]. Skin [PHM^{+}21, SLP21]. Skinning [KB18, WU23]. Smart [XXA^{+}23]. Smith [TE24]. snow [AÖG24]. Soft [CO18, SLP21, XJZ^{+}22]. Soft-Tissue [CO18]. Software [KKT^{+}18, SKW22]. Soil [WFM21]. Soil-Structure [WFM21]. Solids [KBFF^{+}21]. Solver [AHGB19, NZT19]. Solvers [TB23]. solving [GCKM24]. Sonification [LH23]. Sounds [Rüs21]. Southern [Zho24]. Space [Kur24, LHK^{+}24, TY18, PW23, SBE22]. Sparse [HHCM21, RWY^{+}23, STK20]. Spatial [PJL21, TLTM18, Tok23, LH23]. Spatial-Temporal [PJL21]. Spatio [DWBV24]. Spatio-Temporal [DWBV24]. Spatiotemporal [OCW22]. spatiotemporally [AÖG24]. Speak [Far21]. Specialization [SFPO22]. Speech [FLS^{+}22, SG22]. Speech-Driven [FLS^{+}22]. speech-to-image [SG22]. Speed [SNL23]. SPH [HRZ^{+}19, KBFF^{+}21, WJB23]. SPH-centric [HRZ^{+}19]. Sphere [YCL^{+}19].

Spherical [PD19]. Spiking [PPAM23]. Spindle [Kur24]. Spinning [MD21]. Spiral [KWK22]. Splatting [CL24, FHSS18, PKK^{+}24, SGH18]. Spline [ZWWL23]. Split [LVY^{+}20]. Splitting [EK18]. Spoken [MXZ22]. Sprouting [SLMM22]. Stable [TLP^{+}22, XO21]. Stage [GCY^{+}22, Two22]. standard [EHZ23]. Starvation [RN24a]. Starvation-Less [RN24a]. Stereo [SBE22]. Stereo-consistent [SBE22]. Stereoscopic [CNI^{+}20]. Stiff [XXA^{+}23]. Stochastic [DHB24, LY20, PWSF24, VSE21]. Stochastic-Depth [VSE21]. Stories [Zho24]. Strategies [LHR^{+}23]. Stress [SWF^{+}22]. Strike [Oh23]. Stroke [LU24, LZY21]. Strokes [TTC24]. Strong [XXA^{+}23]. Structure [LSL^{+}24, RHH^{+}22, WFM21]. Structure-Preserving [LSL^{+}24]. Study [SLW^{+}21]. Style [MSK22, PJL21, SCNW19, TTK^{+}21]. Stylization [EZ22, PJL21]. Sub [GBW20, TKC21]. Sub-Pixel [TKC21]. Sub-triangle [GBW20]. Subaltern [Far21]. Subdivision [LMSS18]. Subject [MFKK24]. Subspace [YH23]. Subsurface [Hes22, WSD24, XO21, XOKN20]. Succinct [WD24]. Super [WXCT19]. Super-Resolution [WXCT19]. Supersampling [TLP^{+}22]. Supporting [SFPO22]. Surface [CWZ^{+}18, NF24, TWS^{+}24, YCL^{+}19]. Surfaces [HHM23, LMSS18, STSK20, WS18]. SYNC [RW21]. Synchronization [WSG19]. Synplant [HCK23]. Synthesis [CNI^{+}20, PPAM23, PLRD21, STSK20, LSG24, SG22]. Synthesizing [JWB^{+}21]. synthetic [ZCL24]. System [BYH23, FXX^{+}22, Sim21, TTC24, WH21, EHZ23]. Systemic [Cha24]. Systems [NBD^{+}24].

Tailored [FSFS23]. Tale [GCY^{+}22]. Taoist [HCR^{+}21]. Task [FHT^{+}24, GCKM24]. Techniques [OBC22]. Technologies [MD21]. Technology [HCR^{+}21, HK24]. Telepresence [HGR^{+}22]. Temporal [DWBV24, PJL21, SPD18]. Temporally [TLP^{+}22, XO21]. Termination [KDSD24]. Testing [Yuk18]. Text [FLS^{+}22, TWT^{+}23]. Text-to-Gesture [TWT^{+}23]. Textiles [MDD^{+}22]. Texture [PWSF24, STSK20, WJG^{+}21]. Texturing [BD22]. Their [MFKK24]. Thin [WMZ^{+}20]. Three [JDZD19, QLZF21, STSK20, Two22]. Three-Level [STSK20]. Three-party [JDZD19]. Thyself [OSM^{+}22]. Tightly [SWP^{+}23]. Tile [BWL18]. Tile-based [BWL18]. Tiled [FHSS18]. Tiles [MYS20]. Tiling [Kir18, LSG24]. Time [BBGB20, CKY^{+}22, CYH^{+}24, CLS^{+}21, CAS22, DWBV24, GSNK24, KSK24, KOF^{+}24, LY19, LY20, MUEM22, MSK22, PD19, SKW22, TWS^{+}24, TK23, TLP^{+}22, WSD24, XJZ^{+}22, YCL^{+}19, CNI^{+}20, KKSM18, LD18, LPW20, SPD18, SCNW19, TTK^{+}21, WH21, XOKN20, XO21, ZCL18]. Time-independent [BBGB20]. Tissue [CO18, SLP21]. Too [XXA^{+}23]. Toolkit [FG24]. Topologies [BD22, BBM24]. Topology [BWL18]. TouchAR [LMP22]. Traced [XJZ^{+}22]. Tracing [HMES20, HHCM21, KB24, KSK24, LMSS18, MKVH24, WP22, WSD24, GBW20]. Tracking [BBD24, BDWG21, KWK22, NBD^{+}24, SWF^{+}22]. Tracking-Based [SWF^{+}22]. Trading [KDSD24]. Tradition [MD21]. Traditional [HCR^{+}21, MD21, OBC22, EZ24]. trained [PPAM23]. Training [CKY^{+}22]. Trajectory [DMH^{+}21]. Transfer [BDBS22, LWM19, SCNW19, TTK^{+}21, Two22, MMMC21]. Transformations [MSK22, WU23]. Transformed [KHDN22]. Transforming [DHB24]. Transforms [WMZ^{+}20]. Translation [LL23, SLY20]. Transparency [MKKP18]. transparent [GBW20]. Tree [EK18]. Trees [AS22, BD24, Dup20, LVY^{+}20, QLZF21]. Trevor [Day21]. triangle [GBW20]. Trous [DHD24]. Tutorial [CYH^{+}24]. Two [GCY^{+}22, MWW18, NF24, SSH^{+}21, WU23, Yuk24]. Two-Level [NF24]. Two-Player [SSH^{+}21, Yuk24]. Two-Way [MWW18, WU23]. Types [WMZ^{+}20].

Understanding [KKI^{+}18]. Unified [CDGB19, SFPO22, SK23]. Unpaired [LL23]. Unveiling [SPT24]. UpFlow [RPP21]. Using [BBD24, CG23, JWB^{+}21, KBM23, MBN^{+}24, NF24, OBC22, Tok23, WMZ^{+}20, YBB^{+}21, BYH23, DAI^{+}18, ECM23, EZ24, FSFS23, GHH24, HN18, LH23, LSG24, PW23, PPAM23, RPHD20, SGH18, Tri24, WDM24, YW19, YEK24]. Utilizing [HOKY23, TKC21].

Validated [DMH^{+}21]. validation [RDBC24]. Vanished [HK24]. Variance [KDSD24, OCW22, XOKN20]. Variance-Guided [OCW22, XOKN20]. Variates [XO21]. Varied [LSL^{+}24]. Various [FJdNJ22]. Versatile [SLX^{+}23]. Vertex [KKT^{+}18, KKI^{+}18, PKM22]. Vertex-Blend [PKM22]. via [FHT^{+}24, MSK22, PJL21, STK20, SLY20]. Victory [Hes22]. Videos [TTK^{+}21]. View [CNI^{+}20, DAI^{+}18, LBR^{+}18, LBG18, Oh23, PLRD21]. View-Dependent [LBG18]. View-Region [LBR^{+}18]. Vignette [Cha24]. VIPER [ARM^{+}19]. Virtual [BBD24, CYH^{+}24, DPM24, ERHR22, GCKM24, HMN18, LWLY22, MCC22, OSM^{+}22, SLY20, SLW^{+}21, SWF^{+}22, WSX^{+}22, WH21, CASR21]. virtuo [ZM23]. Virus [HHYG24]. Viscous [AHGB19]. Visibility [HMN18, KW21, YEK24]. Vision [BDWG21, JFS^{+}21, Sim21]. Vision-Based [JFS^{+}21]. Visual [ANEK21, CKY^{+}22, RN24b, WFM21]. Visualization [HCK23, KWK22, LZY21, WDG^{+}19, XBN^{+}23]. Visually [PPAM23, YZK^{+}19]. VNDF [TE24]. Voice [GCKM24]. volcanic [LRC^{+}22]. Volume [ARM^{+}19, BDTD22, BMB^{+}24, SLP21, SKRS24, Tri24]. Volumes [HHCM21, HHM23]. Volumetric [BYH23, DAI^{+}18, FG24, HMES20, MLD^{+}24, RCH^{+}23, RN24a, Wan18]. Voxel [vdLSE20]. VPLs [SGH18]. VR [FXX^{+}22, GCY^{+}22, WGD^{+}23]. VR-Based [WGD^{+}23].

Walking [SHS^{+}21]. Walkthrough [RN24a]. Warp [KKS^{+}20]. Waterdrop [Fuj21]. Wavelet [DHD24]. Wavelet-Based [DHD24]. Way [MWW18, WU23]. Ways [BYH23]. Weaving [MD21]. Weights [PHM^{+}21]. Wheelchair [XBN^{+}23]. Wigan [LMP22]. Wind [Zho24]. Windblown [RDBC24]. Wise [MSK22, HOKY23]. Without [DB23]. Woods [SSH^{+}21]. Work [Day21, KOF^{+}24]. Worlds [RN24a]. Woven [MDD^{+}22]. Written [MXZ22, Zho24].

Year [Two22].

Zero [KBFF^{+}21, Kur24, MBN^{+}24]. Zero-Energy [KBFF^{+}21]. Zero-Shot [MBN^{+}24]. ZH3 [RSS^{+}24]. Zonal [RSS^{+}24].

References

Aranjuelo:2024:LGA

[AHAC^{+}24] Nerea Aranjuelo, Siyu Huang, Ignacio Arganda-Carreras, Luis Unzueta, Oihana Otaegui, Hanspeter Pfister, and Donglai Wei. Learning gaze-aware compositional GAN from limited annotations. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(2):28:1–28:??, May 2024. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3654706.

Aanjaneya:2019:EGM

[AHGB19] Mridul Aanjaneya, Chengguizi Han, Ryan Goldade, and Christopher Batty. An efficient geometric multigrid solver for viscous liquids. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(2):14:1–14:21, July 2019. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3340255.

Andersson:2020:FDE

[ANAM^{+}20] Pontus Andersson, Jim Nilsson, Tomas Akenine-Möller, Magnus Oskarsson, Kalle Åström, and Mark D. Fairchild. FLIP: a difference evaluator for alternating images. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(2): 15:1–15:23, August 2020. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3406183.

Andrews:2021:CFV

[ANEK21] Sheldon Andrews, Loic Nassif, Kenny Erleben, and Paul G. Kry. Coupling friction with visual appearance. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(3):31:1–31:20, September 2021. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3480138.

Andreasson:2024:PSA

[AÖG24] Simon Andreasson, Linus Östergaard, and Prashant Goswami. Parallel spatiotemporally adaptive DEM-based snow simulation. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 7(3):50:1–50:??, August 2024. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3675374.

Angles:2019:VVI

[ARM^{+}19] Baptiste Angles, Daniel Rebain, Miles Macklin, Brian Wyvill, Loic Barthe, Jp Lewis, Javier Von Der Pahlen, Shahram Izadi, Julien Valentin, Sofien Bouaziz, and Andrea Tagliasacchi. VIPER: Volume invariant position-based elastic rods. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(2): 19:1–19:26, July 2019. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3340260.

Adams:2022:BFP

[AS22] Andrew Adams and Dillon Sharlet. Better fixed-point filtering with averaging trees. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(3): 32:1–32:??, July 2022. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3543869.

Barkevich:2024:UDL

[BBD24] Kevin Barkevich, Reynold Bailey, and Gabriel J. Diaz. Using deep learning to increase eye-tracking robustness, accuracy, and precision in virtual reality. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(2):27:1–27:??, May 2024. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3654705.

Brunel:2020:TID

[BBGB20] Camille Brunel, Pierre Bénard, Gaël Guennebaud, and Pascal Barla. A time-independent deformer for elastic-rigid contacts. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(1):9:1–9:21, April 2020. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3384539.

Barczak:2024:DDH

[BBM24] Joshua Barczak, Carsten Benthin, and David McAllister. DGF: a dense, hardware-friendly geometry format for lossily compressing meshlets with arbitrary topologies. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(3): 46:1–46:??, August 2024. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3675383.

Barbier:2022:HPH

[BD22] Wilhem Barbier and Jonathan Dupuy. Htex: Per-halfedge texturing for arbitrary mesh topologies. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(3):29:1–29:??, July 2022. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3543868.

Benyoub:2024:CBT

[BD24] Anis Benyoub and Jonathan Dupuy. Concurrent binary trees for large-scale game components. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 7(3):31:1–31:??, August 2024. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3675371.

Belcour:2022:DDP

[BDBS22] Laurent Belcour, Thomas Deliot, Wilhem Barbier, and Cyril Soler. A data-driven paradigm for precomputed radiance transfer. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(3):26:1–26:??, July 2022. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3543864.

Benthin:2022:PPL

[BDTD22] Carsten Benthin, Radoslaw Drabinski, Lorenzo Tessari, and Addis Dittebrandt. PLOC++: Parallel locally-ordered clustering for bounding volume hierarchy construction revisited. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(3): 31:1–31:??, July 2022. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3543867.

Bugdayci:2021:IVE

[BDWG21] Irem Bugdayci, Anne-Heloise Dautel, Robert Wuss, and Ruairi Glynn. Instruments of vision: Eye-tracking and robotics as an embodied interface. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2): 21:1–21:10, July 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3465618.

Bruggisser:2024:RCA

[BLLJ^{+}24] Florian Christoph Bruggisser, Chris Elvis Leisi, Pascal Lund-Jensen, Martin Fröhlich, and Christopher Lloyd Salter. recon-FIGURE: Confronting audiences with digital Doppelgängers. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(4): 52:1–52:??, July 2024. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3664208.

Benthin:2024:HPH

[BMB^{+}24] Carsten Benthin, Daniel Meister, Joshua Barczak, Rohan Mehalwal, John Tsakok, and Andrew Kensler. H-PLOC: Hierarchical parallel locally-ordered clustering for bounding volume hierarchy construction. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(3): 30:1–30:??, August 2024. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3675377.

Bian:2018:TBP

[BWL18] Xiaojun Bian, Li-Yi Wei, and Sylvain Lefebvre. Tile-based pattern design with topology control. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 1(1):23:1–23:15, July 2018. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3203204.

Ban:2023:ISN

[BYH23] Seonghoon Ban, Taeha Yi, and Kyung Hoon Hyun. Intersection of seeing: New ways of experiencing reality using autonomous volumetric capture system. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(2): 19:1–19:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3597632.

Currius:2022:RTH

[CAS22] Roc R. Currius, Ulf Assarsson, and Erik Sintorn. Realtime hair filtering with convolutional neural networks. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1): 15:1–15:15, May 2022. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522606.

Chemistruck:2021:EAP

[CASR21] Mike Chemistruck, Andrew Allen, John Snyder, and Nikunj Raghuvanshi. Efficient acoustic perception for virtual AI agents. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(3):43:1–43:13, September 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3480139.

Chang:2019:USM

[CDGB19] Jumyung Chang, Fang Da, Eitan Grinspun, and Christopher Batty. A unified simplicial model for mixed-dimensional and nonmanifold deformable elastic objects. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(2):11:1–11:18, July 2019. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3340252.

Cole:2023:KCU

[CG23] Adam Cole and Mick Grierson. Kiss/crash: Using diffusion models to explore real desire in the shadow of artificial representations. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(2): 17:1–17:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3597625.

Chan:2024:SRV

[Cha24] Ka Chi Chan. ’Sisyphus’: a robotic vignette of systemic conflict and democratic endeavors. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(4):54:1–54:??, July 2024. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3664220.

Cardoso:2022:TPV

[CKY^{+}22] Joao Liborio Cardoso, Bernhard Kerbl, Lei Yang, Yury Uralsky, and Michael Wimmer. Training and predicting visual error for real-time applications. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1): 11:1–11:17, May 2022. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3522625.

Chen:2024:DGG

[CL24] Wenbo Chen and Ligang Liu. Deblur-GS: 3D Gaussian splatting from camera motion blurred images. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 7(1):18:1–18:??, May 2024. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3651301.

Chermain:2021:RTG

[CLS^{+}21] Xavier Chermain, Simon Lucas, Basile Sauvage, Jean-Michel Dischler, and Carsten Dachsbacher. Real-time geometric glint antialiasing with normal map filtering. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1):1:1–1:16, April 2021. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3451257.

Chaurasia:2020:PRT

[CNI^{+}20] Gaurav Chaurasia, Arthur Nieuwoudt, Alexandru-Eugen Ichim, Richard Szeliski, and Alexander Sorkine-Hornung. Passthrough+: Real-time stereoscopic view synthesis for mobile mixed reality. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(1):7:1– 7:17, April 2020. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3384540.

Casas:2018:LNS

[CO18] Dan Casas and Miguel A.

Otaduy. Learning nonlinear soft-tissue dynamics for interactive avatars. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1):10:1–10:15, July 2018. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3203187.

Chen:2018:R

[CSN18] Ge Chen, Pedro V. Sander, and Diego Nehab. The replate. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1): 4:1–4:14, July 2018. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3203205.

Chen:2018:DSL

[CWZ^{+}18] Anpei Chen, Minye Wu, Yingliang Zhang, Nianyi Li, Jie Lu, Shenghua Gao, and Jingyi Yu. Deep surface light fields. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1): 14:1–14:17, July 2018. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3203192.

Chen:2024:ITM

[CYH^{+}24] Boyuan Chen, Xinan Yan, Xuning Hu, Dominic Kao, and Hai-Ning Liang. Impact of tutorial modes with different time flow rates in virtual reality games. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(1):6:1–6:??, May 2024. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3651296.

Cao:2023:IEC

[CYS23] Chong Cao, Lingfei Yang, and Xukun Shen. Immersive experience of Chinese poems and paintings. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(2): 18:1–18:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3597621.

Delanoy:2018:SUM

[DAI^{+}18] Johanna Delanoy, Mathieu Aubry, Phillip Isola, Alexei A. Efros, and Adrien Bousseau. 3D sketching using multi-view deep volumetric prediction. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1): 21:1–21:22, July 2018. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3203197.

Day:2021:RDP

[Day21] Kevin Day. Reversal, disconnect, and proposition: Noise and data politics in the work of Julian Oliver and Trevor Paglen. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2): 25:1–25:8, July 2021. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3465624.

Ding:2023:DCN

[DB23] Xinwen Ding and Christopher Batty. Differentiable curl-noise: Boundary-respecting procedural incompressible flows without discontinuities. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(1):16:1–16:??, May 2023. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3585511.

Denisova:2024:CAA

[DB24] Elena Denisova and Leonardo Bocchi. Converging algorithmagnostic denoising for Monte Carlo rendering. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(3): 40:1–40:??, August 2024. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3675384.

DiBlasi:2023:BEM

[DCB23] Johnny DiBlasi, Carlos Castellanos, and Bello Bello. Beauty: Explorations of machine-microbial agencies. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(2):25:1–25:??, August 2023. CODEN ???? ISSN 2577-6193 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/10.1145/3597626.

deDinechin:2022:RLM

[dDB22] Heloise de Dinechin and Laurent Belcour. Rendering layered materials with diffuse interfaces. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1):13:1–13:12, May 2022. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522620.

Deliot:2024:TND

[DHB24] Thomas Deliot, Eric Heitz, and Laurent Belcour. Transforming a non-differentiable rasterizer into a differentiable one with stochastic gradient estimation. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(1): 3:1–3:??, May 2024. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3651298.

Dolp:2024:FGS

[DHD24] Reiner Dolp, Johannes Hanika, and Carsten Dachsbacher. A fast GPU schedule for À-trous wavelet-based denoisers. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(1): 15:1–15:??, May 2024. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3651299.

Daniel:2021:PVM

[DMH^{+}21] Beatríz Cabrero Daniel, Ricardo Marques, Ludovic Hoyet, Julien Pettré, and Josep Blat. A perceptually-validated metric for crowd trajectory quality evaluation. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(3):42:1–42:18, September 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3480136.

Dinkov:2024:PHL

[DPM24] Martin Dinkov, Sumanta Pattanaik, and Ryan P. McMahan. Perceptions of hybrid lighting for virtual reality. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(1): 5:1–5:??, May 2024. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3651292.

Dupuy:2020:CBT

[Dup20] Jonathan Dupuy. Concurrent binary trees (with application to longest edge bisection). Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(2): 21:1–21:20, August 2020. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3406186.

Donnelly:2024:FFA

[DWBV24] William Donnelly, Alan Wolfe, Judith Bütepage, and Jon Valdés. FAST: Filter-adapted spatio-temporal sampling for real-time rendering. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(1): 13:1–13:??, May 2024. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3651283.

Dong:2021:EAR

[DXS21] Yihong Dong, Lunchen Xie, and Qingjiang Shi. Efficient algorithms for rotation averaging problems. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1):16:1–16:16, April 2021. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3451263.

Egan:2023:NQE

[ECM23] Dónal Egan, Darren Cosker, and Rachel McDonnell. NeuroDog: Quadruped embodiment using neural networks. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(3): 38:1–38:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3606936.

Elran:2023:FMS

[EHZ23] Sharan R. Elran, Yuval Harel, and Amit R. Zoran. Fine motor skills: Operating standard robotic fabrication as a generative system. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(2):

30:1–30:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3597627.

Estevez:2018:ISM

[EK18] Alejandro Conty Estevez and Christopher Kulla. Importance sampling of many lights with adaptive tree splitting. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(2): 25:1–25:17, August 2018. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3233305.

Eckert:2022:PLR

[ERHR22] Marie Eckert, Thomas Robotham, Emanuël A. P. Habets, and Olli S. Rummukainen. Pupillary light reflex correction for robust pupillometry in virtual reality. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(2):18:1–18:??, May 2022. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3530798.

Elran:2022:GMD

[EZ22] Sharan R. Elran and Amit R. Zoran. The ghost in the machine: Digital stylization and expressive re-materialization of ancient ceramics. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(4):44:1–44:??, September 2022. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3533609.

Elran:2024:PCM

[EZ24] Sharan R. Elran and Amit R.

Zoran. Probabilistic craft- materialization of generated images using digital and traditional craft. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(4):63:1–63:??, July 2024. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3664209.

Farahi:2021:CSS

[Far21] Behnaz Farahi. Can the subaltern speak?: Feminism in robotic mask design. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2): 18:1–18:11, July 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3465621.

Fox-Gieg:2024:LAT

[FG24] Nick Fox-Gieg. Lightning artist toolkit: a hand-drawn volumetric animation pipeline. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(4): 58:1–58:??, July 2024. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3664221.

Franke:2018:MLD

[FHSS18] Linus Franke, Nikolai Hofmann, Marc Stamminger, and Kai Selgrad. Multi-layer depth of field rendering with tiled splatting. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1):6:1–6:17, July 2018. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3203200.

Fan:2024:ADR

[FHT^{+}24] Shu-Ho Fan, Kai-Wen Hsiao, Kai Yi Tan, Chih-Yuan Yao, and Hung-Kuo Chu. Aliasing detection in rendered images via a multi-task learning. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(3): 41:1–41:??, August 2024. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3675379.

Faria:2022:IPV

[FJdNJ22] Matheus Prado Prandini Faria, Etienne Silva Julia, Marcelo Zanchetta do Nascimento, and Rita Maria Silva Julia. Investigating the performance of various deep neural networks-based approaches designed to identify game events in gameplay footage. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1):8:1–8:17, May 2022. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3522624.

Fan:2022:JAT

[FLS^{+}22] Yingruo Fan, Zhaojiang Lin, Jun Saito, Wenping Wang, and Taku Komura. Joint audiotext model for expressive speechdriven 3D facial animation. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1): 16:1–16:15, May 2022. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522615.

Fink:2023:ERL

[FSFS23] Laura Fink, Svenja Strobel, Linus Franke, and Marc Stamminger. Efficient rendering for light field displays using tailored projective mappings. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(1): 3:1–3:??, May 2023. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3585498.

Fujioka:2021:DIA

[Fuj21] Sadam Fujioka. drop: an interactive art installation with waterdrop projection-mapping. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2):27:1–27:8, July 2021. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3465613.

Fu:2022:EEC

[FXX^{+}22] Zhiying Fu, Rui Xu, Shiqing Xin, Shuangmin Chen, Changhe Tu, Chenglei Yang, and Lin Lu. EasyVRModeling: Easily create 3D models by an immersive VR system. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1): 10:1–10:14, May 2022. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522613.

Gruen:2020:STO

[GBW20] Holger Gruen, Carsten Benthin, and Sven Woop. Sub-triangle opacity masks for faster ray tracing of transparent objects. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(2): 18:1–18:12, August 2020. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3406180.

Guo:2024:CMD

[GCKM24] Siqi Guo, Minsoo Choi, Dominic Kao, and Christos Mousas. Collaborating with my Doppelgänger: The effects of selfsimilar appearance and voice of a virtual character during a jigsaw puzzle co-solving task. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(1): 4:1–4:??, May 2024. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3651288.

Gochfeld:2022:TTP

[GCY^{+}22] David Gochfeld, Alex Coulombe, Yu-Jun Yeh, Robert Lester, Robert Barry Fleming, Zachary Meicher-Buzzi, and Ari Tarr. A tale of two productions: a Christmas Carol on stage and in VR. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(4):41:1–41:??, September 2022. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3533612.

Granizo-Hidalgo:2024:IRC

[GHH24] Ana Granizo-Hidalgo and Nicolas Holzschuch. Interactive rendering of caustics using dimension reduction for manifold next-event estimation. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(1): 12:1–12:??, May 2024. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3651297.

Goel:2024:RTD

[GSNK24] Rahul Goel, Markus Schütz, P. J. Narayanan, and Bernhard Kerbl. Real-time decompression and rasterization of massive point clouds. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(3): 48:1–48:??, August 2024. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3675373.

Gavane:2023:EBP

[GW23] Ajinkya Gavane and Benjamin Watson. Eye-based point rendering for dynamic multiview effects. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(1):5:1–5:??, May 2023. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3585513.

Gavane:2024:LFD

[GW24] Ajinkya Gavane and Benjamin Watson. Light field display point rendering. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(1):19:1–19:??, May 2024. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3651300.

Hu:2023:SCV

[HCK23] Youyang Hu, Chiaochi Chou, and Yasuaki Kakehi. Synplant: Cymatics visualization of plant-environment interaction based on plants biosignals. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(2): 26:1–26:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3597622.

Han:2018:DB

[HCNS18] Songfang Han, Ge Chen, Diego Nehab, and Pedro V. Sander. In-depth buffers. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1):2:1–2:14, July 2018. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3203194.

Henchoz:2021:MSD

[HCR^{+}21] Nicolas Henchoz, Margaux Charvolin, Delphine Ribes, Lara Défayes, Cédric DuchÊne, Emily Groves, and Andreas Sonderegger. Ming Shan digital experience: Immersive technology for traditional Taoist meditation. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2):24:1–24:10, July 2021. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3465620.

Hessels:2022:BVS

[Hes22] Scott Hessels. Below victory:

Subsurface radar advances for creative digital heritage. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5 (4):47:1–47:??, September 2022. CODEN ???? ISSN 2577-6193 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3533389.

Hildreth:2019:CMA

[HG19] Dalton Hildreth and Stephen J.

Guy. Coordinating multi-agent navigation by learning communication. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 2(2):20:1–20:17, July 2019. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3340261.

Han:2019:HMP

[HGG^{+}19] Xuchen Han, Theodore F. Gast, Qi Guo, Stephanie Wang, Chenfanfu Jiang, and Joseph Teran. A hybrid material point method for frictional contact with diverse materials. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(2):17:1–17:24, July 2019. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3340258.

Hossaini:2022:GTA

[HGR^{+}22] Ali Hossaini, Oliver Gingrich, Shama Rahman, Mick Grierson, Joshua Murr, Alan Chamberlain, and Alain Renaud. GROUPTHINK: Telepresence and agency during live performance. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(4):39:1–39:??, September 2022. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3533610.

Hofmann:2021:IPT

[HHCM21] Nikolai Hofmann, Jon Hasselgren, Petrik Clarberg, and Jacob Munkberg. Interactive path tracing and reconstruction of sparse volumes. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1): 5:1–5:19, April 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3451256.

Hofmann:2023:JND

[HHM23] Nikolai Hofmann, Jon Hasselgren, and Jacob Munkberg. Joint neural denoising of surfaces and volumes. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(1): 10:1–10:??, May 2023. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3585497.

Huang:2024:ELV

[HHYG24] Jiayang Huang, Yue Huang, David Yip, and Varvara Guljajeva. Ephemera: Language as a virus — AI-driven interactive and immersive art installation. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(4):62:1–62:??, July 2024. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3664219.

Hui:2024:CVC

[HK24] Tak-Cheung Hui and Yu-Chia Kuo. A concert in a vanished church: Contextualizing Peace Island’s auditory history with modern technology. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(4): 65:1–65:??, July 2024. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3664218.

Habermann:2023:HHA

[HLX^{+}23] Marc Habermann, Lingjie Liu, Weipeng Xu, Gerard Pons-Moll, Michael Zollhoefer, and Christian Theobalt. HDHumans: a hybrid approach for highfidelity digital humans. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(3): 36:1–36:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3606927.

Hofmann:2020:NDP

[HMES20] Nikolai Hofmann, Jana Martschinke, Klaus Engel, and Marc Stamminger. Neural denoising for path tracing of medical volumetric data. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 3(2):13:1–13:18, August 2020. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3406181.

Hunt:2018:HVV

[HMN18] Warren Hunt, Michael Mara, and Alex Nankervis. Hierarchical visibility for virtual reality. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1):8:1–8:18, July 2018. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3203191.

Heitz:2018:HPE

[HN18] Eric Heitz and Fabrice Neyret.

High-performance by-example noise using a histogram-preserving blending operator. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(2): 31:1–31:25, August 2018. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3233304.

Han:2023:PWG

[HOKY23] Kyu Beom Han, Olivia G. Odenthal, Woo Jae Kim, and Sung-Eui Yoon. Pixel-wise guidance for utilizing auxiliary features in Monte Carlo denoising. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(1): 11:1–11:??, May 2023. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3585505.

Holzer:2022:SPI

[Hol22] Derek Holzer. In search of the plastic image: a media archaeology of scan processing. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5 (4):45:1–45:??, September 2022. CODEN ???? ISSN 2577-6193 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3539218.

Huang:2019:GNP

[HRZ^{+}19] Kemeng Huang, Jiming Ruan, Zipeng Zhao, Chen Li, Changbo Wang, and Hong Qin. A general novel parallel framework for SPH-centric algorithms. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(1): 7:1–7:16, June 2019. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3321360.

Huang:2023:TRH

[HYW^{+}23] Li Huang, Fan Yang, Chendi Wei, Yu Ju (Edwin) Chen, Chun Yuan, and Ming Gao. Towards realtime: a hybrid physics-based method for hair animation on GPU. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(3): 43:1–43:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3606937.

Jin:2019:DLB

[JDZD19] Aobo Jin, Qixin Deng, Yuting Zhang, and Zhigang Deng. A deep learning-based model for head and eye motion generation in three-party conversations. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(2):9:1–9:19, July 2019. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3340250.

Jonnalagadda:2021:RVB

[JFS^{+}21] Aditya Jonnalagadda, Iuri Frosio, Seth Schneider, Morgan McGuire, and Joohwan Kim. Robust vision-based cheat detection in competitive gaming. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1): 7:1–7:18, April 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3451259.

Jambon:2023:NIE

[JKK^{+}23] Clément Jambon, Bernhard Kerbl, Georgios Kopanas, Stavros Diolatzis, Thomas Leimkühler, and George Drettakis. NeRFshop: Interactive editing of neural radiance fields. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(1): 1:1–1:??, May 2023. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3585499.

Jiang:2021:SIS

[JWB^{+}21] Hao Jiang, Siqi Wang, Huikun Bi, Xiaolei Lv, Binqiang Zhao, Zheng Wang, and Zhaoqi Wang. Synthesizing indoor scene layouts in complicated architecture using dynamic convolution networks. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 4(1):13:1–13:16, April 2021. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3451267.

Jia:2023:PCA

[JWLC23] Shiyang Jia, Stephanie Wang, Tzu-Mao Li, and Albert Chern. Physical cyclic animations. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(3): 45:1–45:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3606938.

Komaritzan:2018:PS

[KB18] Martin Komaritzan and Mario Botsch. Projective skinning. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1):12:1–12:19, July 2018. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3203203.

Kacerik:2024:SOK

[KB24] Martin Kácerik and Jirí Bittner. SAH-optimized k-DOP hierarchies for ray tracing. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(3): 32:1–32:??, August 2024. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3675391.

Kugelstadt:2021:FCE

[KBFF^{+}21] Tassilo Kugelstadt, Jan Bender, José Antonio Fernández-Fernández, Stefan Rhys Jeske, Fabian Löschner, and Andreas Longva. Fast corotated elastic SPH solids with implicit zero-energy mode control. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4 (3):33:1–33:21, September 2021. CODEN ???? ISSN 2577-6193 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3480142.

Krogmeier:2023:UEF

[KBM23] Claudia Krogmeier, Esteban García Bravo, and Christos Mousas. Using experimental filmmaking to create an engaging braincomputer interface. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(2): 29:1–29:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3597624.

Kandlbinder:2024:OPT

[KDSD24] Lukas Kandlbinder, Addis Dittebrandt, Alexander Schipek, and Carsten Dachsbacher. Optimizing path termination for radiance caching through explicit variance trading. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(3): 33:1–33:??, August 2024. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3675381.

KT:2022:BLT

[KHDN22] Aakash KT, Eric Heitz, Jonathan Dupuy, and P. J. Narayanan.

Bringing linearly transformed cosines to anisotropic GGX. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1):12:1–12:18, May 2022. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3522612.

Kirillov:2018:NPT

[Kir18] Aleksandr Kirillov. Non-periodic tiling of procedural noise functions. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 1(2):32:1–32:15, August 2018. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3233306.

Kaur:2022:RMB

[KJM22] Harsimran Kaur, Swati Jindal, and Roberto Manduchi. Rethinking model-based gaze estimation. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 5(2):17:1–17:??, May 2022. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3530797.

Kadlecek:2019:BAP

[KK19] Petr Kadlecek and Ladislav Kavan. Building accurate physicsbased face models from data. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(2):15:1–15:16, July 2019. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3340256.

Kerbl:2018:RVC

[KKI^{+}18] Bernhard Kerbl, Michael Kenzel, Elena Ivanchenko, Dieter Schmalstieg, and Markus Steinberger. Revisiting the vertex cache: Understanding and optimizing vertex processing on the modern GPU. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(2): 29:1–29:16, August 2018. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3233302.

Kim:2020:PRW

[KKS^{+}20] Joohwan Kim, Pyarelal Knowles, Josef Spjut, Ben Boudaoud, and Morgan Mcguire. Postrender warp with late input sampling improves aiming under high latency conditions. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(2): 12:1–12:18, August 2020. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3406187.

Koniaris:2018:GAD

[KKSM18] Babis Koniaris, Maggie Kosek, David Sinclair, and Kenny Mitchell. GPU-accelerated depth codec for real-time, highquality light field reconstruction. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1):3:1–3:15, July 2018. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3203193.

Kenzel:2018:FVR

[KKT^{+}18] Michael Kenzel, Bernhard Kerbl, Wolfgang Tatzgern, Elena Ivanchenko, Dieter Schmalstieg, and Markus Steinberger. On-the-fly vertex reuse for massively-parallel software geometry processing. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(2): 28:1–28:17, August 2018. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3233303.

Kim:2021:FMO

[KLXvdP21] Nam Hee Kim, Hung Yu Ling, Zhaoming Xie, and Michiel van de Panne. Flexible motion optimization with modulated assistive forces. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(3):35:1–35:25, September 2021. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3480144.

Kuth:2024:RTP

[KOF^{+}24] Bastian Kuth, Max Oberberger, Carsten Faber, Dominik Baumeister, Matthäus Chajdas, and Quirin Meyer. Realtime procedural generation with GPU work graphs. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(3): 47:1–47:??, August 2024. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3675376.

Kelling:2024:LPG

[KSK24] Jan Kelling, Daniel Ströter, and Arjan Kuijper. Light path guided culling for hybrid realtime path tracing. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(3): 37:1–37:??, August 2024. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3675387.

Kurbak:2024:HZG

[Kur24] Ebru Kurbak. Handcrafting in zero gravity: Reinventing the spindle as an artistic intervention in space research. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(4): 64:1–64:??, July 2024. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3664216.

Koch:2021:GVS

[KW21] Thomas Koch and Michael Wimmer. Guided visibility sampling++. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1):4:1–4:16, April 2021. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3451266.

Koch:2022:SMG

[KWK22] Maurice Koch, Daniel Weiskopf, and Kuno Kurzhals. A spiral into the mind: Gaze spiral visualization for mobile eye tracking. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(2):20:1–20:??, May 2022. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3530795.

Lachenmyer:2022:AMP

[LA22] Nathan S. Lachenmyer and Sadiya Akasha. An aquarium of machines: a physically realized artificial life simulation. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5 (4):34:1–34:??, September 2022. CODEN ???? ISSN 2577-6193 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3533388.

Li:2018:ERB

[LAJB18] Xiaokai Li, Sheldon Andrews, Ben Jones, and Adam Bargteil. Energized rigid body fracture. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1):9:1–9:9, July 2018. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3203207.

Li:2019:MRM

[LB19] Yijing Li and Jernej Barbic.

Multi-resolution modeling of shapes in contact. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(2):12:1–12:26, July 2019. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3340253.

Lambert:2018:VDM

[LBG18] Thibaud Lambert, Pierre Bénard, and Gaël Guennebaud. A viewdependent metric for patchbased LOD generation 8 selection. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1):20:1–20:21, July 2018. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3203195.

Lall:2018:VRO

[LBR^{+}18] Puneet Lall, Silviu Borac, Dave Richardson, Matt Pharr, and Manfred Ernst. Viewregion optimized image-based scene simplification. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(2): 26:1–26:22, August 2018. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3233311.

Lee:2018:RTA

[LD18] Haebom Lee and Piotr Didyk.

Real-time apparent resolution enhancement for head-mounted displays. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1):19:1–19:15, July 2018. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3203202.

Loschner:2023:MEP

[LFFJ^{+}23] Fabian Löschner, José Antonio Fernández-Fernández, Stefan Rhys Jeske, Andreas Longva, and Jan Bender. Micropolar elasticity in physically-based animation. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 6(3):46:1–46:??, August 2023. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3606922.

Leischner:2023:SJP

[LH23] Vojtech Leischner and Pavel Husa. Sonification of a juggling performance using spatial audio. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 6(2):20:1–20:??, August 2023. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3597619.

Lin:2024:CSA

[LHK^{+}24] Rem RunGu Lin, Botao Amber Hu, Koo Yongen Ke, Wei Wu, and Kang Zhang. Cell space: Augmented awareness of intercorporeality. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(4): 53:1–53:??, July 2024. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3664213.

Liu:2024:SPI

[LHL24] Zitan Liu, Yikai Huang, and Ligang Liu. ShaderPerFormer: Platform-independent contextaware shader performance predictor. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 7(1):2:1–2:??, May 2024. CODEN ???? ISSN 2577-6193 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3651295.

Liu:2023:IBR

[LHR^{+}23] Zihao Liu, Jing Huang, Allan Rocha, Jim Malmros, and Jerry Zhang. Importance-based ray strategies for dynamic diffuse global illumination. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(1): 9:1–9:??, May 2023. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3585500.

Lee:2024:PDM

[LJC24] Fei Lee, Jia-Wun Jhang, and Chun-Fa Chang. Photon-driven manifold sampling. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(3):

34:1–34:??, August 2024. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3675375.

Li:2023:UTP

[LL23] Chih-Chia Li and I-Chen Lin. Unpaired translation of 3D point clouds with multi-part shape representation. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(1):6:1–6:??, May 2023. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3585508.

Lewis:2022:ERW

[LMP22] Chara Lewis, Kristin Mojsiewicz, and Anneke Pettican. Erratics on the road to Wigan Pier: The creation of TouchAR. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(4):40:1–40:??, September 2022. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3533611.

Lier:2018:HRC

[LMSS18] Alexander Lier, Magdalena Martinek, Marc Stamminger, and Kai Selgrad. A highresolution compression scheme for ray tracing subdivision surfaces with displacement. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(2): 33:1–33:17, August 2018. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3233308.

Luksch:2020:RTA

[LPW20] Christian Luksch, Lukas Prost, and Michael Wimmer. Realtime approximation of photometric polygonal lights. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(1): 4:1–4:18, April 2020. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3384537.

Lastic:2022:ISP

[LRC^{+}22] Maud Lastic, Damien Rohmer, Guillaume Cordonnier, Claude Jaupart, Fabrice Neyret, and Marie-Paule Cani. Interactive simulation of plume and pyroclastic volcanic ejections. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1): 4:1–4:15, May 2022. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522609.

Lutz:2024:FOA

[LSG24] Nicolas Lutz, Arnaud Schoentgen, and Guillaume Gilet. Fast orientable aperiodic ocean synthesis using tiling and blending. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 7(3):49:1–49:??, August 2024. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3675388.

Li:2024:EDL

[LSL^{+}24] Tianxing Li, Rui Shi, Zihui Li, Takashi Kanai, and Qing Zhu. Efficient deformation learning of varied garments with a structure-preserving multilevel framework. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(1):8:1–8:??, May 2024. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3651286.

Levien:2024:GFS

[LU24] Raph Levien and Arman Uguray. GPU-friendly stroke expansion. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 7(3):35:1–35:??, August 2024. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3675390.

Lin:2020:HAD

[LVY^{+}20] Daqi Lin, Elena Vasiou, Cem Yuksel, Daniel Kopta, and Erik Brunvand. Hardwareaccelerated dual-split trees. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(2): 20:1–20:21, August 2020. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3406185.

Li:2022:VBN

[LWLY22] Tianyu Li, Wenyou Wang, Daqi Lin, and Cem Yuksel. Virtual blue noise lighting. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(3): 23:1–23:??, July 2022. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3543872.

Li:2019:DPR

[LWM19] Yue Li, Pablo Wiedemann, and Kenny Mitchell. Deep precomputed radiance transfer for deformable objects. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(1):3:1– 3:16, June 2019. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3320284.

Lin:2019:RTR

[LY19] Daqi Lin and Cem Yuksel.

Real-time rendering with lighting grid hierarchy. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(1):8:1– 8:17, June 2019. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3321361.

Lin:2020:RTS

[LY20] Daqi Lin and Cem Yuksel. Realtime stochastic lightcuts. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(1): 5:1–5:18, April 2020. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3384543.

Lyu:2021:ISM

[LZY21] Ruimin Lyu, Tianqin Zhang, and Zhaolin Yuan. Imaginary stroke movement measurement and visualization. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2): 20:1–20:12, July 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3465625.

Mercier-Aubin:2023:ARD

[MAK23] Alexandre Mercier-Aubin and Paul G. Kry. Adaptive rigidification of discrete shells. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(3): 39:1–39:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3606932.

Maquiling:2024:ZSS

[MBN^{+}24] Virmarie Maquiling, Sean Anthony Byrne, Diederick C. Niehorster, Marcus Nyström, and Enkelejda Kasneci. Zeroshot segmentation of eye features using the segment anything model (SAM). Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(2):26:1–26:??, May 2024. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3654704.

Mandal:2022:IPB

[MCC22] Avirup Mandal, Parag Chaudhuri, and Subhasis Chaudhuri. Interactive physics-based virtual sculpting with haptic feedback. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1):9:1–9:20, May 2022. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522611.

Montero:2021:SCI

[MD21] Valentina Montero and Sandra De Berduccy. Spinning the conductors of an indigenous tradition: Toward a continuity of traditional Andean weaving with new electronic technologies. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2): 26:1–26:12, July 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3465622.

Meiklejohn:2022:WBO

[MDD^{+}22] Elizabeth Meiklejohn, Felicita Devlin, John Dunnigan, Patricia Johnson, Joy Xiaoji Zhang, Steve Marschner, Brooks Hagan, and Joy Ko. Woven behavior and ornamentation: Simulation-assisted design and application of self-shaping woven textiles. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(4):37:1–37:??, September 2022. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3533682.

Meng:2018:KFR

[MDZV18] Xiaoxu Meng, Ruofei Du, Matthias Zwicker, and Amitabh Varshney. Kernel foveated rendering. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 1(1):5:1–5:20, July 2018. CODEN ???? ISSN 2577-6193 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3203199.

Macklin:2020:LOR

[MEM^{+}20] Miles Macklin, Kenny Erleben, Matthias Müller, Nuttapong Chentanez, Stefan Jeschke, and Zach Corse. Local optimization for robust signed distance field collision. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(1):8:1–8:17, April 2020. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3384538.

Meshi:2024:GMH

[Mes24] Avital Meshi. GPT-ME: a human–AI cognitive assemblage. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(4):55:1–55:??, July 2024. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3664214.

Melnyk:2024:PSO

[MFKK24] Kateryna Melnyk, Lee Friedman, Dmytro Katrychuk, and Oleg Komogortsev. Per-subject oculomotor plant mathematical models and the reliability of their parameters. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(2): 24:1–24:??, May 2024. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3654701.

Munstermann:2018:MBO

[MKKP18] Cedrick Münstermann, Stefan Krumpen, Reinhard Klein, and Christoph Peters. Momentbased order-independent transparency. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1):7:1–7:20, July 2018. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3203206.

Meister:2024:HRT

[MKVH24] Daniel Meister, Paritosh Kulkarni, Aaryaman Vasishta, and Takahiro Harada. HIPRT: a ray tracing framework in HIP. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(3): 44:1–44:??, August 2024. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3675378.

Medin:2024:FMR

[MLD^{+}24] Safa C. Medin, Gengyan Li, Ruofei Du, Stephan Garbin, Philip Davidson, Gregory W. Wornell, Thabo Beeler, and Abhimitra Meka. FaceFolds: Meshed radiance manifolds for efficient volumetric rendering of dynamic faces. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(1):23:1–23:??, May 2024. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3651304.

Musoni:2021:FST

[MMMC21] Pietro Musoni, Riccardo Marin, Simone Melzi, and Umberto Castellani. A functional skeleton transfer. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(3):34:1–34:15, September 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3480140.

Mason:2022:RTS

[MSK22] Ian Mason, Sebastian Starke, and Taku Komura. Realtime style modelling of human locomotion via featurewise transformations and local motion phases. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1):

6:1–6:18, May 2022. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522618.

Macek:2022:RTR

[MUEM22] Nejc Macek, Baran Usta, Elmar Eisemann, and Ricardo Marroquim. Real-time relighting of human faces with a low-cost setup. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1):14:1–14:19, May 2022. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522626.

Mukai:2018:SBR

[Muk18] Tomohiko Mukai. Samplingbased rig conversion into nonrigid helper bones. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1): 13:1–13:17, July 2018. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3203190.

Mukherjee:2018:ITW

[MWW18] Rajaditya Mukherjee, Longhua Wu, and Huamin Wang. Interactive two-way shape design of elastic bodies. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1): 11:1–11:17, July 2018. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3203196.

Maher:2022:FCW

[MXZ22] Kevin Maher, Fan Xiang, and Liang Zhi. FaceType: Crafting written impressions of spoken expression. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(4):38:1–38:??, September 2022. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3533385.

Mallett:2020:EAD

[MYS20] Ian Mallett, Cem Yuksel, and Larry Seiler. Efficient adaptive deferred shading with hardware scatter tiles. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(2): 11:1–11:17, August 2020. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3406184.

Nguyen:2024:DDA

[NBD^{+}24] Viet Dung Nguyen, Reynold Bailey, Gabriel J. Diaz, Chengyi Ma, Alexander Fix, and Alexander Ororbia. Deep domain adaptation: a Sim2Real neural approach for improving eyetracking systems. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(2): 25:1–25:??, May 2024. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3654703.

Nikolic:2023:ARA

[NBM23] Predrag K. Nikoli’c, Giacomo Bertin, and Kishan Munroe. Aquaterrestrial recolonization: AI environmentalist toward design for climate action related behaviour change. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(2): 24:1–24:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3597630.

Niehaus:2021:MBA

[NF21] Kiona Hagen Niehaus and Rebecca Fiebrink. Making up 3D bodies: Artistic and serendipitous modeling of digital human figures. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2):23:1–23:9, July 2021. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3468779.

Nishidate:2024:EPB

[NF24] Yuki Nishidate and Issei Fujishiro. Efficient particle-based fluid surface reconstruction using mesh shaders and bidirectional two-level grids. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(1): 1:1–1:??, May 2024. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3651285.

Narain:2019:SOA

[NZT19] Rahul Narain, Jonas Zehnder, and Bernhard Thomaszewski. A second-order advection-reflection solver. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 2(2):16:1–16:14, July 2019. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3340257.

Odefunso:2022:TAD

[OBC22] Adebunmi E. Odefunso, Esteban Garcia Bravo, and Yingjie V. Chen. Traditional African dances preservation using deep learning techniques. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5 (4):35:1–35:??, September 2022. CODEN ???? ISSN 2577-6193 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3533608.

Oberberger:2022:SVG

[OCW22] Max Oberberger, Matthäus G. Chajdas, and Rüdiger Westermann. Spatiotemporal varianceguided filtering for motion blur. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(3):22:1–22:??, July 2022. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3543871.

Ogaki:2020:GLP

[Oga20] Shinji Ogaki. Generalized light portals. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 3(2):10:1–10:19, August 2020. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3406176.

Oh:2023:KDR

[Oh23]  Jooyoung Oh.  The kestrel

drone: Reimagining bird’s eye view with biomimetic AI-drone for bird strike prevention. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(2): 27:1–27:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3597629.

Oliver:2022:DMD

[OSM^{+}22] Marilène Oliver, Scott Smallwood, Stephan Moore, J. R. Carpenter, and Jonathan Cohn. Dissecting my data body: How to know thyself as a virtual reality in the digital age. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(4):42:1–42:??, September 2022. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3533387.

Peters:2019:SPS

[PD19] Christoph Peters and Carsten Dachsbacher. Sampling projected spherical caps in real time. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(1):1:1–1:16, June 2019. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3320282.

Pan:2021:HHN

[PHM^{+}21] Xiaoyu Pan, Jiancong Huang, Jiaming Mai, He Wang, Honglin Li, Tongkui Su, Wenjun Wang, and Xiaogang Jin. HeterSkinNet: a heterogeneous network for skin weights prediction. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1): 10:1–10:19, April 2021. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3451262.

Park:2021:DMS

[PJL21] Soomin Park, Deok-Kyeong Jang, and Sung-Hee Lee. Diverse motion stylization for multiple style domains via spatialtemporal graph-based generative model. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(3):36:1–36:17, September 2021. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3480145.

Papantonakis:2024:RMF

[PKK^{+}24] Panagiotis Papantonakis, Georgios Kopanas, Bernhard Kerbl, Alexandre Lanvin, and George Drettakis. Reducing the memory footprint of 3D Gaussian splatting. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(1):16:1–16:??, May 2024. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3651282.

Peters:2022:PCV

[PKM22] Christoph Peters, Bastian Kuth, and Quirin Meyer. Permutation coding for vertex-blend attribute compression. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1): 5:1–5:16, May 2022. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522607.

Prakash:2021:HIB

[PLRD21] Siddhant Prakash, Thomas Leimkühler, Simon Rodriguez, and George Drettakis. Hybrid image-based rendering for freeview synthesis. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1):15:1–15:20, April 2021. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3451260.

Polykretis:2023:IFV

[PPAM23] Ioannis Polykretis, Aditi Patil, Mridul Aanjaneya, and Konstantinos Michmizos. An interactive framework for visually realistic 3D motion synthesis using evolutionarily-trained spiking neural networks. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(1): 12:1–12:??, May 2023. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3585509.

Papatheodorou:2023:LDF

[PW23] Theodoros Papatheodorou and Jessica Wolpert. Lights! Dance! Freeze!: Exploring the dancemusical filmic space using embodied search in an interactive installation. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(2): 21:1–21:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3597620.

Pharr:2024:FAS

[PWSF24] Matt Pharr, Bartlomiej Wronski, Marco Salvi, and Marcos Fajardo. Filtering after shading with stochastic texture filtering. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(1):14:1–14:??, May 2024. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3651293.

Quigley:2021:TDR

[QLZF21] Ed Quigley, Winnie Lin, Yilin Zhu, and Ronald Fedkiw. Three dimensional reconstruction of botanical trees with simulatable geometry. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(3):37:1–37:16, September 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3480146.

Razon:2023:LAM

[RCH^{+}23] Alan Marquez Razon, Yizhou Chen, Yushan Han, Steven Gagniere, Michael Tupek, and Joseph Teran. A linear and angular momentum conserving hybrid particle/grid iteration for volumetric elastic contact. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(3): 44:1–44:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3606924.

Rosset:2024:WSA

[RDBC24] Nicolas Rosset, Regis Duvigneau, Adrien Bousseau, and Guillaume Cordonnier. Windblown sand around obstacles — simulation and validation of deposition patterns. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(1): 20:1–20:??, May 2024. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3651284.

Reshetov:2022:RRI

[Res22]  Alexander Reshetov.  Ray/

ribbon intersections. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(3):

28:1–28:??, July 2022. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3543862.

Raateland:2022:DAG

[RHH^{+}22] Wouter Raateland, Torsten Hädrich, Jorge Alejandro Amador Herrera, Daniel T. Banuti, Wojciech Pałubicki, Sören Pirk, Klaus Hildebrandt, and Dominik L. Michels. DCGrid: an adaptive grid structure for memory-constrained fluid simulation on the GPU. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1): 3:1–3:14, May 2022. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3522608.

Reshetov:2018:PRH

[RL18] Alexander Reshetov and David Luebke. Phantom ray-hair intersector. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 1(2):34:1–34:22, August 2018. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3233307.

Richermoz:2024:GDS

[RN24a] Antoine Richermoz and Fabrice Neyret. GigaVoxels DP: Starvation-less render and production for large and detailed volumetric worlds walkthrough. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 7(3):42:1–42:??, August 2024. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3675389.

Robillard:2024:CCM

[RN24b] Gaëtan Robillard and Jérôme Nika. Critical climate machine: a visual and musical exploration of climate misinformation through machine learning. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(4):56:1–56:??, July 2024. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3664215.

Rodriguez:2020:IBR

[RPHD20] Simon Rodriguez, Siddhant Prakash, Peter Hedman, and George Drettakis. Image-based rendering of cars using semantic labels and approximate reflection flow. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(1):6:1–6:17, April 2020. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3384535.

Roy:2021:NUS

[RPP21] Bruno Roy, Pierre Poulin, and Eric Paquette. Neural Up-Flow: a scene flow learning approach to increase the apparent resolution of particlebased liquids. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(3):40:1–40:26, September 2021. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3480147.

Roughton:2024:ZQZ

[RSS^{+}24] Thomas Roughton, Peter-Pike Sloan, Ari Silvennoinen, Michal Iwanicki, and Peter Shirley. ZH3: Quadratic zonal harmonics. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(1):11:1–11:??, May 2024. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3651294.

Rust:2021:BMG

[Rüs21] Annina Rüst. Bad mother /good mother: The poetics and politics of the sounds of invisible labor. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2):17:1–17:9, July 2021. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3465617.

Riahi:2021:PBN

[RW21] Maryam Riahi and Benjamin Allen Watson. Am I playing better now?: The effects of G-SYNC in 60Hz gameplay. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1): 6:1–6:17, April 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3451269.

Reda:2023:PBM

[RWY^{+}23] Daniele Reda, Jungdam Won, Yuting Ye, Michiel van de Panne, and Alexander Winkler. Physics-based motion retargeting from sparse inputs. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(3): 33:1–33:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3606928.

Shi:2022:SCS

[SBE22] Peiteng Shi, Markus Billeter, and Elmar Eisemann. Stereoconsistent screen-space ambient occlusion. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1):2:1–2:12, May 2022. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3522614.

Smith:2019:ENN

[SCNW19] Harrison Jesse Smith, Chen Cao, Michael Neff, and Yingying Wang. Efficient neural networks for real-time motion style transfer. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(2):13:1–13:17, July 2019. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3340254.

Su:2018:ISB

[SDY^{+}18] Wanchao Su, Dong Du, Xin Yang, Shizhe Zhou, and Hongbo Fu. Interactive sketch-based normal map generation with deep neural networks. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1): 22:1–22:17, July 2018. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3203186.

Seitz:2022:SUS

[SFPO22] Kerry A. Seitz, Theresa Foley, Serban D. Porumbescu, and John D. Owens. Supporting unified shader specialization by co-opting C++ features. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(3): 25:1–25:??, July 2022. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3543866.

Sola:2022:DPE

[SG22] Mar Canet Sola and Varvara Guljajeva. Dream painter: Exploring creative possibilities of AI-aided speech-to-image synthesis in the interactive art context. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(4):33:1–33:??, September 2022. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3533386.

Sriwasansak:2018:EEC

[SGH18] Jamorn Sriwasansak, Adrien Gruson, and Toshiya Hachisuka. Efficient energy-compensated VPLs using photon splatting. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1):16:1–16:13, July 2018. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3203189.

Schmelter:2021:IBR

[SHS^{+}21] Thereza Schmelter, Levente Hernadi, Marc Aurel Störmer, Frank Steinicke, and Kristian Hildebrand. Interaction based redirected walking. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1):9:1–9:16, April 2021. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3451264.

Schutz:2024:SSL

[SHW24] Markus Schütz, Lukas Herzberger, and Michael Wimmer. SimLOD: Simultaneous LOD generation and rendering for point clouds. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(1):17:1–17:??, May 2024. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3651287.

Simon:2021:JVE

[Sim21] Gilles Simon. Jan van Eyck’s perspectival system elucidated through computer vision. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2): 22:1–22:8, July 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3465623.

Shi:2023:UAP

[SK23] Alvin Shi and Theodore Kim. A unified analysis of penaltybased collision energies. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(3): 41:1–41:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3606934.

Steiner:2024:FVC

[SKRS24] Michael Steiner, Thomas Köhler, Lukas Radl, and Markus Steinberger. Frustum volume caching for accelerated NeRF rendering. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 7(3):39:1–39:??, August 2024. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3675370.

Schutz:2022:SRB

[SKW22] Markus Schütz, Bernhard Kerbl, and Michael Wimmer. Software rasterization of 2 billion points in real time. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(3):

24:1–24:??, July 2022. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3543863.

Silva:2022:FLI

[SLMM22] Pedro Silva, Daniel Lopes, Pedro Martins, and Penousal Machado. Field of leaves: an interactive installation sprouting leaves out of public money. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(4):43:1–43:??, September 2022. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3533683.

Sheen:2021:VPS

[SLP21] Seung Heon Sheen, Egor Larionov, and Dinesh K. Pai. Volume preserving simulation of soft tissue with skin. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4 (3):32:1–32:23, September 2021. CODEN ???? ISSN 2577-6193 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3480143.

Sethapakdi:2022:KIC

[SLR^{+}22] Ticha Sethapakdi, Mackenzie Leake, Catalina Monsalve Rodriguez, Miranda J. Cai, and Stefanie Mueller. KineCAM: an instant camera for animated photographs. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(4):46:1–46:??, September 2022. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3533613.

Shi:2021:VRS

[SLW^{+}21] Rongkai Shi, Hai-Ning Liang, Yu Wu, Difeng Yu, and Wenge Xu. Virtual reality sickness mitigation methods: a comparative study in a racing game. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1): 8:1–8:16, April 2021. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3451255.

Su:2023:GCM

[SLX^{+}23] Haozhe Su, Xuan Li, Tao Xue, Chenfanfu Jiang, and Mridul Aanjaneya. A generalized constitutive model for versatile MPM simulation and inverse learning with differentiable physics. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(3): 49:1–49:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3606925.

Seiler:2020:CCG

[SLY20] Larry Seiler, Daqi Lin, and Cem Yuksel. Compacted CPU/GPU data compression via modified virtual address translation. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(2): 19:1–19:18, August 2020. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3406177.

Sin:2023:NPA

[SNL23] Zackary P. T. Sin, Peter H. F.

Ng, and Hong Va Leong. NeR-Fahedron: a primitive for animatable neural rendering with interactive speed. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(1): 2:1–2:??, May 2023. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3585512.

Schied:2018:GER

[SPD18] Christoph Schied, Christoph Peters, and Carsten Dachsbacher. Gradient estimation for realtime adaptive temporal filtering. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 1(2):24:1–24:16, August 2018. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3233301.

Schreiner:2021:GPP

[SPL^{+}21] Paul Schreiner, Maksym Perepichka, Hayden Lewis, Sune Darkner, Paul G. Kry, Kenny Erleben, and Victor B. Zordan. Global position prediction for interactive motion capture. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4 (3):39:1–39:16, September 2021.

CODEN ???? ISSN 2577-6193 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3479985.

Sobhan:2024:UNA

[SPT24] Arshia Sobhan, Philippe Pasquier, and Adam Tindale. Unveiling new artistic dimensions in calligraphic Arabic script with generative adversarial networks. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(4): 61:1–61:??, July 2024. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3664210.

Swearingen:2021:WMR

[SSH^{+}21] Kyoung Swearingen, Scott Swearingen, Fede Camara Halac, Sruthi Ammannagari, and Matt Hall. The Woods: a mixed-reality two-player cooperative game. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2):30:1–30:7, July 2021. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3465616.

Starke:2023:MBP

[SSKS23] Paul Starke, Sebastian Starke, Taku Komura, and Frank Steinicke. Motion in-betweening with phase manifolds. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(3): 37:1–37:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3606921.

Schuster:2020:HPI

[STK20] Kersten Schuster, Philip Trettner, and Leif Kobbelt. Highperformance image filters via sparse approximations. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(2): 14:1–14:19, August 2020. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3406182.

Schuster:2020:TLA

[STSK20] Kersten Schuster, Philip Trettner, Patric Schmitz, and Leif Kobbelt. A three-level approach to texture mapping and synthesis on 3D surfaces. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(1): 1:1–1:19, April 2020. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3384542.

Stoeve:2022:ETB

[SWF^{+}22] Maike Stoeve, Markus Wirth, Rosanna Farlock, André Antunovic, Victoria Müller, and Bjoern M. Eskofier. Eye tracking-based stress classification of athletes in virtual reality. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(2):19:1–19:??, May 2022. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3530796.

Shi:2023:LCM

[SWP^{+}23] Alvin Shi, Haomiao Wu, Jarred Parr, A. M. Darke, and Theodore Kim. Lifted curls: a model for tightly coiled hair simulation. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 6(3):42:1–42:??, August 2023. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3606920.

Stuyck:2023:DDP

[SyC23] Tuur Stuyck and Hsiao yu Chen. DiffXPBD: Differentiable positionbased simulation of compliant constraint dynamics. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(3): 51:1–51:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3606923.

Takahashi:2023:MAS

[TB23] Tetsuya Takahashi and Christopher Batty. A multilevel activeset preconditioner for boxconstrained pressure Poisson solvers. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 6(3):50:1–50:??, August 2023. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3606939.

Tokuyoshi:2024:BVS

[TE24] Yusuke Tokuyoshi and Kenta Eto. Bounded VNDF sampling for the Smith–GGX BRDF. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(1): 10:1–10:??, May 2024. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3651291.

Temkin:2024:OPL

[Tem24] Daniel Temkin. The Olympus programming language. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(4): 60:1–60:??, July 2024. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3664211.

Taylor:2023:RTS

[TK23] Brennen Taylor and John Keyser. Real-time sand dune simulation. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(1):15:1–15:??, May 2023. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3585510.

Tumanov:2021:DDP

[TKC21] Evgenii Tumanov, Dmitry Korobchenko, and Nuttapong Chentanez. Data-driven particlebased liquid simulation with deep learning utilizing sub-pixel convolution. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1):12:1–12:16, April 2021. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3451261.

Thomas:2022:TSR

[TLP^{+}22] Manu Mathew Thomas, Gabor Liktor, Christoph Peters, Sungye Kim, Karthik Vaidyanathan, and Angus G. Forbes. Temporally stable realtime joint neural denoising and supersampling. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(3):21:1–21:??, July 2022. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3543870.

Tang:2018:PPS

[TLTM18] Min Tang, Zhongyuan Liu, Ruofeng Tong, and Dinesh Manocha. PSCC: Parallel selfcollision culling with spatial hashing on GPUs. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1):18:1–18:18, July 2018. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3203188.

Tokuyoshi:2023:ESR

[Tok23] Yusuke Tokuyoshi. Efficient spatial resampling using the PDF similarity. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(1):4:1–4:??, May 2023. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3585501.

Tricard:2024:ISU

[Tri24] Thibault Tricard. Interval shading: using mesh shaders to generate shading intervals for volume rendering. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(3): 43:1–43:??, August 2024. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3675380.

Tong:2024:ESE

[TTC24] Derek C. W. Tong, Xuet Ying Tan, and Alex Q. Chen. Eye strokes: an eye-gaze drawing system for Mandarin characters. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(2):29:1–29:??, May 2024. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3654702.

Texler:2021:FIR

[TTK^{+}21] Aneta Texler, Ondrej Texler, Michal Kucera, Menglei Chai, and Daniel Sýkora. FaceBlit: Instant real-time example-based style transfer to facial videos. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1):14:1–14:17, April 2021. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3451270.

Twomey:2022:TSD

[Two22] Robert Twomey. Three stage drawing transfer: Collaborative drawing between a generative adversarial network, corobotic arm, and five-year-old child. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(4):36:1–36:??, September 2022. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3533614.

Twomey:2023:CCA

[Two23] Robert Twomey. Communing with creative AI. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(2): 28:1–28:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3597633.

Tatzgern:2024:RCS

[TWS^{+}24] Wolfgang Tatzgern, Alexander Weinrauch, Pascal Stadlbauer, Joerg H. Mueller, Martin Winter, and Markus Steinberger. Radiance caching with on-surface caches for real-time global illumination. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(3): 38:1–38:??, August 2024. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3675382.

Teshima:2023:AAB

[TWT^{+}23] Hitoshi Teshima, Naoki Wake, Diego Thomas, Yuta Nakashima, Hiroshi Kawasaki, and Katsushi Ikeuchi. ACT2G: Attentionbased contrastive learning for text-to-gesture generation. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(3): 35:1–35:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3606940.

Truong:2018:NRF

[TY18] Nghia Truong and Cem Yuksel.

A narrow-range filter for screenspace fluid rendering. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1): 17:1–17:15, July 2018. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3203201.

Truong:2020:QAC

[TYS20] Nghia Truong, Cem Yuksel, and Larry Seiler. Quadratic approximation of cubic curves. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(2): 16:1–16:17, August 2020. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3406178.

Talukdar:2024:LCM

[TZW24] Bilas Talukdar, Yunhao Zhang, and Tomer Weiss. Learning crowd motion dynamics with crowds. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(1):9:1–9:??, May 2024. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3651302.

vanderLaan:2020:LGC

[vdLSE20] Remi van der Laan, Leonardo Scandolo, and Elmar Eisemann. Lossy geometry compression for high resolution voxel scenes. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(1):2:1–2:13, April 2020. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3384541.

Viitanen:2018:PFH

[VKJ^{+}18] Timo Viitanen, Matias Koskela, Pekka Jääskeläinen, Aleksi Tervo, and Jarmo Takala. PLOCTree: a fast, high-quality hardware BVH builder. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(2): 35:1–35:19, August 2018. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3233309.

Vermeer:2021:SDA

[VSE21] Jop Vermeer, Leonardo Scandolo, and Elmar Eisemann.

Stochastic-depth ambient occlusion. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1):3:1–3:15, April 2021. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3451268.

Wang:2018:FMC

[Wan18] Cuilan Wang. A fast method to calculate the volumetric divergence metric for evaluating the accuracy of the extracted isosurface. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(2): 27:1–27:19, August 2018. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3233312.

Waldemarson:2024:SOM

[WD24] Gustaf Waldemarson and Michael Doggett. Succinct opacity micromaps. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(3):45:1–45:??, August 2024. CODEN ???? ISSN 2577-6193 (electronic).

URL  https://dl.acm.org/
doi/10.1145/3675385.

Wang:2019:SVD

[WDG^{+}19] Stephanie Wang, Mengyuan Ding, Theodore F. Gast, Leyi Zhu, Steven Gagniere, Chenfanfu Jiang, and Joseph M. Teran. Simulation and visualization of ductile fracture with the material point method. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(2): 18:1–18:20, July 2019. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3340259.

Wang:2024:PFL

[WDM24] Binglun Wang, Niladri Shekhar Dutt, and Niloy J. Mitra. ProteusNeRF: Fast lightweight NeRF editing using 3D-aware image context. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(1):22:1–22:??, May 2024. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3651290.

Weiss:2023:FPB

[Wei23] Tomer Weiss. Fast positionbased multi-agent group dynamics. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(1):14:1–14:??, May 2023. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3585507.

Wang:2021:VSS

[WFM21] Xu Wang, Makoto Fujisawa, and Masahiko Mikawa. Visual simulation of soil-structure destruction with seepage flows. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4 (3):41:1–41:18, September 2021. CODEN ???? ISSN 2577-6193 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3480141.

Wan:2023:MVB

[WGD^{+}23] Dayu Wan, Xiaolei Guo, Jiahui Dong, Christos Mousas, and Yingjie Chen. ManiLoco: a VR-based locomotion method for concurrent object manipulation. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(1):7:1–7:??, May 2023. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3585502.

Wu:2021:MGA

[WH21] Ziwei Wu and Lingdong Huang.

Mimicry: Genetic-algorithm-based real-time system of virtual insects in a living environment — a new and altered nature. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2):28:1–28:8, July 2021. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3465615.

Westhofen:2023:CLC

[WJB23] Lukas Westhofen, Stefan Jeske, and Jan Bender. A comparison of linear consistent correction methods for first-order SPH derivatives. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(3): 48:1–48:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3606933.

Wu:2021:RGI

[WJG^{+}21] Jane Wu, Yongxu Jin, Zhenglin Geng, Hui Zhou, and Ronald Fedkiw. Recovering geometric information with learned texture perturbations. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(3):38:1–38:18, September 2021. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3480137.

Wu:2023:EAB

[WK23] Haomiao Wu and Theodore Kim. An eigenanalysis of anglebased deformation energies. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(3): 40:1–40:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3606929.

Wald:2020:UHR

[WMZ^{+}20] Ingo Wald, Nate Morrical, Stefan Zellmann, Lei Ma, Will Usher, Tiejun Huang, and Valerio Pascucci. Using hardware ray transforms to accelerate ray/ primitive intersections for long, thin primitive types. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(2): 17:1–17:16, August 2020. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3406179.

Wald:2022:DPP

[WP22] Ingo Wald and Steven G. Parker.

Data parallel path tracing with object hierarchies. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(3): 30:1–30:??, July 2022. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3543861.

Wang:2018:FGP

[WS18] Tong Wang and Reiji Suda. Fast generation of Poisson-disk samples on mesh surfaces by progressive sample projection. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(2): 30:1–30:18, August 2018. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3233310.

Werner:2024:RSS

[WSD24] Mirco Werner, Vincent Schüßler, and Carsten Dachsbacher. Re-STIR subsurface scattering for real-time path tracing. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(3): 36:1–36:??, August 2024. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3675372.

Watson:2019:EAS

[WSG19] Benjamin Watson, Rachit Shrivastava, and Ajinkya Gavane. The effects of adaptive synchronization on performance and experience in gameplay. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(1): 5:1–5:13, June 2019. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3320286.

Wang:2022:ERR

[WSX^{+}22] Jialin Wang, Rongkai Shi, Zehui Xiao, Xueying Qin, and Hai-Ning Liang. Effect of render resolution on gameplay experience, performance, and simulator sickness in virtual reality games. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1):7:1–7:15, May 2022. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522610.

Wu:2023:TWC

[WU23] Yuhan Wu and Nobuyuki Umetani. Two-way coupling of skinning transformations and position based dynamics. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(3): 47:1–47:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3606930.

Werhahn:2019:MPG

[WXCT19] Maximilian Werhahn, You Xie, Mengyu Chu, and Nils Thuerey. A multi-pass GAN for fluid flow super-resolution. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(2): 10:1–10:21, July 2019. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3340251.

Xie:2023:MQV

[XBN^{+}23] Yurui Xie, Giulia Barbareschi, Ayesha Nabila, Kai Kunze, and Masa Inakage. Movement quality visualization for wheelchair dance. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 6(2):23:1–23:??, August 2023. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3597628.

Xu:2022:RTR

[XJZ^{+}22] Yang Xu, Yuanfa Jiang, Junbo Zhang, Kang Li, and Guohua Geng. Real-time ray-traced soft shadows of environmental lighting by conical ray culling. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1): 1:1–1:15, May 2022. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522617.

Xu:2021:GLA

[XK21] Pei Xu and Ioannis Karamouzas. A GAN-like approach for physics-based imitation learning and interactive character control. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(3):44:1–44:22, September 2021. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3480148.

Xie:2021:RTS

[XO21] Tiantian Xie and Marc Olano. Real-time subsurface control variates: Temporally stable adaptive sampling. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1):2:1– 2:18, April 2021. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3451265.

Xie:2020:RTS

[XOKN20] Tiantian Xie, Marc Olano, Brian Karis, and Krzysztof Narkowicz. Real-time subsurface scattering with single pass variance-guided adaptive importance sampling. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(1):3:1–3:21, April 2020. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3384536.

Xie:2023:HPC

[XTS^{+}23] Zhaoming Xie, Jonathan Tseng, Sebastian Starke, Michiel van de Panne, and C. Karen Liu. Hierarchical planning and control for box loco-manipulation. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(3): 31:1–31:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3606931.

Xie:2023:TST

[XXA^{+}23] Kaixiang Xie, Pei Xu, Sheldon Andrews, Victor B. Zordan, and Paul G. Kry. Too stiff, too strong, too smart: Evaluating fundamental problems with motion control policies. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(3): 34:1–34:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3606935.

Yi:2021:RUB

[YBB^{+}21] Changyoon Yi, Juhyun Bae, Nakkyu Baek, Jina Jung, Sunwoong Hur, Hyun Jean Lee, and Seung Ah Lee. ReMember: Using biosignals to recall memories of companion animals. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2): 29:1–29:7, July 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3465614.

Yang:2019:RTF

[YCL^{+}19] Bowen Yang, William Corse, Jiecong Lu, Joshuah Wolper, and Chen-Fanfu Jiang. Realtime fluid simulation on the surface of a sphere. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(1):4:1– 4:17, June 2019. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3320285.

Ying:2024:EVA

[YEK24] Zhi Ying, Nicholas Edwards, and Mikhail Kutuzov. Efficient visibility approximation for game AI using neural omnidirectional distance fields. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(1): 21:1–21:??, May 2024. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3651289.

Yoshimura:2023:SCR

[YH23] Atsushi Yoshimura and Takahiro Harada. Subspace culling for ray–box intersection. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(1): 8:1–8:??, May 2023. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3585503.

Yokoyama:2024:FBP

[YIFO24] Takumi Yokoyama, Kazuya Izumi, Tatsuki Fushimi, and Yoichi Ochiai. Floating on the boundary: Perceptions of reality in a half-digital half-physical bunny. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 7(4):57:1–57:??, July 2024. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3664217.

Younes:2023:MMA

[YKK^{+}23] Mohamed Younes, Ewa Kijak, Richard Kulpa, Simon Malinowski, and Franck Multon. MAAIP: Multi-agent adversarial interaction priors for imitation from fighting demonstrations for physics-based characters. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(3): 32:1–32:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3606926.

Yuksel:2018:ADA

[Yuk18] Cem Yuksel. Alpha distribution for alpha testing. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1): 1:1–1:11, July 2018. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3203185.

Yuksel:2022:HPP

[Yuk22] Cem Yuksel. High-performance polynomial root finding for graphics. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(3):27:1–27:??, July 2022. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3543865.

Yuksel:2024:SBM

[Yuk24] Cem Yuksel. Skill-based matchmaking for competitive twoplayer games. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(1):7:1–7:??, May 2024. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3651303.

Yang:2019:ABC

[YW19] Hsiang-Yu Yang and Sai-Keung Wong. Agent-based cooperative animation for box-manipulation using reinforcement learning. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(1):2:1–2:18, June 2019. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3320283.

Yang:2021:EHO

[YY21] Zeshi Yang and Zhiqi Yin. Efficient hyperparameter optimization for physics-based character animation. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1):11:1–11:19, April 2021. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3451254.

Yang:2019:VLC

[YZK^{+}19] Lei Yang, Dmitry Zhdan, Emmett Kilgariff, Eric B. Lum, Yubo Zhang, Matthew Johnson, and Henrik Rydgård. Visually lossless content and motion adaptive shading in games. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(1): 6:1–6:19, June 2019. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3320287.

Zanni:2018:HHR

[ZCL18] Cédric Zanni, Frédéric Claux, and Sylvain Lefebvre. HCSG: Hashing for real-time CSG modeling. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 1(1):15:1–15:19, July 2018. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3203198.

Zhang:2024:RCS

[ZCL24] Weidi Zhang, Lijiaozi Cheng, and Jieliang Luo. ReCollection: Creating synthetic memories with AI in an interactive art installation. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(4): 51:1–51:??, July 2024. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3664207.

Zhou:2024:SWA

[Zho24] Tengchao Zhou. Southern wind algorist: Merging nine algorithms with nine gay stories written in classical Chinese. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 7(4): 59:1–59:??, July 2024. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3664212.

Zeidler:2023:BVS

[ZM23] Daniel Zeidler and Matthew McGinity. Bodylab: in virtuo sculpting, painting and performing of full-body avatars. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(2): 22:1–22:??, August 2023. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3597631.

Zhang:2021:CPI

[ZRL21] Weidi Zhang, Donghao Ren, and George Legrady. Cangjie’s poetry: an interactive art experience of a semantic humanmachine reality. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2):19:1–19:9, July 2021. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3465619.

Zhao:2023:GBB

[ZWWL23] Yuming Zhao, Zhongke Wu, Xingce Wang, and Xinyue Liu.

G2 blending ball B-spline curve by B-spline. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 6(1):13:1–13:??, May 2023. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3585504.